### v1

# Metrics on completely positive maps via noncommutative geometry

Abstract
By considering an infinite-dimensional analogue of the Choi-Jamiolkowski isomorphism, we study how to induce metrics on a distinguished subset of the completely positive maps between tracial $C^{*}$-algebras using seminorms from noncommutative geometry. Under suitable conditions on the these seminorms, we show that the induced metrics will satisfy the quantum information theoretic properties of stability and chaining. Lastly, we show how to generate such metrics from Kasparov exterior products of spectral triples.
footnotetext: Emails: areaus@math.uio.no, bedos@math.uio.no, jonaeid@math.uio.no, nadiasl@math.uio.no, and trono@math.uio.no.
1 Introduction
Trace-preserving completely positive maps are ubiquitous in Quantum Information Theory as, among others, the means to model evolution of quantum systems, these generally being described by matrix algebras. Through a standard duality perspective, there is a companion class of unital, completely positive maps, and these in turn are prominent in Operator Algebras and Noncommutative Geometry, especially in relation to deep questions about approximation properties of operator algebras (in an infinite-dimensional setting). Among many questions which have been asked about trace-preserving completely positive maps one finds those about existence of a metric structure, or distance, capable of capturing permanence properties such as chaining, a form of coherent composition, or stability under tensorial expansion of the system with an ancilla system [BV25, GLN05] present. In this work we propose a framework for developing a metric structure on the set of (unital) completely positive maps between two, not necessarily finite-dimensional $C^{*}$-algebras, by employing constructions native to Noncommutative Geometry.
Attempts so far at developing a metric structure have concentrated on the finite-dimensional setting, namely matrix algebras [BV25], or, in the operator algebraic framework, on the von Neumann algebraic setting, where normal states are a core ingredient [AJL25, AJW25, Daw25].
The main thrust of the present work is that by employing completely positive maps between tracial $C^{*}$-algebras and formalizing a $C^{*}$-algebraic version of an embedding akin to the Choi-Jamiolkowski map in the finite-dimensional setting, we are in a position to take advantage of the constructions of quantum metrics, or distances, on state spaces of $C^{*}$-algebras pioneered by Connes and Rieffel, see e.g.[Con89, Rie98, Rie99, Rie02, Rie04].
Explicitly, we specify a suitable subset of the completely positive maps between $C^{*}$-algebras $A$ and $B$, where $A$ is unital and $B$ is endowed with a faithful trace $\tau_{B}$, which we call trace channels (see Definition 4.13). These are exactly the completely positive maps which yield states under the infinite-dimensional version of the Choi-Jamiolkowski map. We are then able to construct metrics on the space of trace channels by employing techniques from noncommutative geometry, and put forward definitions of stability and chaining in an infinite-dimensional setting. Moreover, we find sufficient conditions for stability and chaining to hold, and illustrate their validity in naturally occuring examples by once more appealing to constructions in noncommutative geometry. Indeed, we show stability for quantum metrics arising from spectral triples (cf. Theorem 6.2), and prove that chaining is satisfied for a natural class of maps between group $C^{*}$-algebras equipped with metrics arising from spectral triples coming from proper length functions on the group, see Theorem 6.9. A large class of examples exhibiting the best possible behavior where chaining and stability hold simultaneously is provided by countable amenable groups with natural length functions, see Example 6.10. The construction of Lipschitz seminorms coming from spectral triples is of importance in the study of compact quantum metric spaces in noncommutative geometry. It has been investigated in a variety of different settings, e.g. for group $C^{*}$-algebras [CR17, OR05], crossed products [AKK25, Haw+13, Kli24], and quantum groups [AK18, AK26, BVZ15, KK25], though this is far from a complete list. In a related line of research, [AJL25] investigates Lipschitz seminorms (which a priori do not come from spectral triples) on quantum permutation groups for the purpose of studying the quantum Wasserstein distance.
Questions such as having stability and chaining are not standard in the context of quantum metrics in noncommutative geometry, and we hope this work opens up a new fruitful line of interaction with quantum information theory.
As mentioned, a large part of the present work relies on formalizing an infinite-dimensional analogue of the Choi-Jamiolkowski map for $C^{*}$-algebras. Related work has appeared in several recent papers: In the quantum graph literature an analogue appears for finite-dimensional $C^{*}$-algebras, cf. [Was24, CGW25]. In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiolkowski isomorphism, see [Haa21, HKS24] and references therein.
The technical difficulties one encounters when moving from a von Neumann algebraic setup to a $C^{*}$-algebraic setting pertain mainly to the need to work with state spaces of tensor products of $C^{*}$-algebras and their opposite algebras. This difficulty does not appear in the finite-dimensional setting as $M_{n}(\mathbb{C})$ is nuclear and isomorphic to its opposite algebra. The latter of these facts is not true for general $C^{*}$-algebras. Thus, without requiring nuclearity of the $C^{*}$-algebras in question, certain maps and their continuity properties very much depend on whether one employs the minimal or the maximal tensor product. A celebrated result of Kirchberg, cf. [Kir93, Kir94], generalizing a landmark result of Connes, cf. [Con76], establishes continuity with respect to the minimal tensor product of a natural map $\mu_{\tau}$, cf. Theorem 2.2. The map $\mu_{\tau}$ is dependent on a trace $\tau$, and Kirchberg’s Theorem states that $\mu_{\tau}$ is continuous in the spatial norm if and only if the corresponding trace $\tau$ is amenable. This allows us to define a tractable version of the Choi-Jamiolkowski map for maps between $C^{*}$-algebras, where the target $C^{*}$-algebra is unital and admits an amenable trace, cf. Theorem 4.15.
The structure of the paper is as follows. We collect a number of important notions and set notation in the preliminary section Section 2. While we focus on a specific way of inducing metrics on certain subsets of completely positive maps in this article, this is not the only induction scheme one could employ. As such, we discuss induction of metrics in Section 3. It also serves to explain the motivation behind the remaining sections. In Section 4 we study the $C^{*}$-algebraic Choi-Jamiolkowski map and use it to introduce the notion of trace channels. The $C^{*}$-algebraic Choi-Jamiolkowski map is then used in Section 5 to induce metrics on the set of trace channels between $C^{*}$-algebras. We especially focus on criteria guaranteeing that these metrics satisfy stability and chaining. Lastly, in Section 6 we use machinery from noncommutative geometry to show that the necessary conditions for stability and chaining from Section 5 are automatically satisfied for metrics arising from the external Kasparov product of spectral triples. We in particular focus on spectral triples for group $C^{*}$-algebras arising from proper length functions on countable groups.
Acknowledgments
This research was funded by The Research Council of Norway [project 324944] and [project 345433].
2 Preliminaries
2.1 Tensor norms, opposite algebras, and traces
By a *-algebra, we will mean a complex algebra $A$ equipped with an involutive, anti-multiplicative, conjugate-linear map, usually denoted by $a\mapsto a^{*}$ for $a\in A$. An isomorphism between *-algebras will then mean an algebra isomorphism which intertwines the involutions. A norm $\|\cdot\|$ on a *-algebra AA is called a $C^{*}$-norm when it is submultiplicative, invariant under the involution and satisfies the $C^{*}$-equality $\|a^{*}a\|=\|a\|^{2}$ for all $a\in A$. The completion of $A$ with respect to a $C^{*}$-norm on $A$ becomes a $C^{*}$-algebra after extending the multiplication and the involution by continuity.
If $A$ and $B$ are $C^{*}$-algebras, then the algebraic tensor product $A\odot B$ of $A$ and $B$ becomes a *-algebra with respect to the involution determined by $(a\otimes b)^{*}=a^{*}\otimes b^{*}$ for all $a\in A$ and $b\in B$ (see for instance [Mur90], p. 188-189). Moreover, if $\delta$ is a $C^{*}$-norm on $A\odot B$, then we will denote the completion of $A\odot B$ with respect to $\delta$ by $A\otimes_{\delta}B$. There are two $C^{*}$-norms in particular that are natural to consider.
First: The spatial norm $\|\cdot\|_{\rm min}$ on $A\odot B$ is defined for any $x\in A\odot B$ by
$\|x\|_{\rm min}=\|(\pi_{A}\otimes\pi_{B})(x)\|_{B(H_{A}\otimes H_{B})},$
where $\pi_{A}\colon A\to B(H_{A})$ is any faithful *-representation of $A$, and $\pi_{B}\colon B\to B(H_{B})$ is any faithful *-representation of $B$. It is well-known that the spatial norm is independent of the choices of $\pi_{A}$ and $\pi_{B}$. The spatial norm satisfies $\|x\|_{\rm min}\leq\delta(x)$ for any $x\in A\odot B$ and any $C^{*}$-norm $\delta$ on $A\odot B$. This is known as Takesaki’s Theorem cf. [Mur90, Section 6.4] or [BO08, Section 3.4] for a proof. We denote the completion of $A\odot B$ with respect to the spatial norm by $A\otimes_{\rm min}B$ and call this the minimal tensor product of $A$ and $B$.
Second: The maximal norm $\|\cdot\|_{\rm max}$ on $A\odot B$ is defined for any $x\in A\odot B$ by
$\|x\|_{\rm max}=\sup\{\|\pi(x)\|_{B(H)}:\pi\colon A\odot B\to B(H)\text{ is a cyclic $*$-representation}\}.$
We denote the completion of $A\odot B$ with respect to the maximal norm by $A\otimes_{\rm max}B$ and call this the maximal tensor product of $A$ and $B$. The maximal tensor product satisfies the following universal property: If $\pi\colon A\odot B\to C$ is any *-homomorphism, where $C$ is a $C^{*}$-algebra, then there exists a unique *-homomorphism $A\otimes_{\rm max}B\to C$ extending $\pi$, cf. [BO08, Proposition 3.3.7].
The universality of the maximal tensor product implies the following: If $\delta$ is a $C^{*}$-norm on $A\odot B$, then there exists a unique *-homomorphism $q_{\delta}\colon A\otimes_{\rm max}B\to A\otimes_{\delta}B$ making the following diagram commute
${A\otimes_{\rm max}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q_{\delta}}$${A\odot B\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{\rm max}}$$\scriptstyle{i_{\delta}}$${A\otimes_{\delta}B,}$
where $i_{\rm max}$ and $i_{\delta}$ are the respective inclusions of $A\odot B$. By continuity of all the *-homomorphisms involved, and density of $A\odot B$, we get that $q_{\delta}$ is surjective for any $C^{*}$-norm $\delta$. A $C^{*}$-algebra $A$ is nuclear if $q_{\delta}$ is injective for all $C^{*}$-algebras $B$ and any $C^{*}$-norm $\delta$ on $A\odot B$. By Takesaki’s Theorem, we can equivalently say that $A$ is nuclear if $A\otimes_{\rm max}B$ is *-isomorphic to $A\otimes_{\rm min}B$ for every $C^{*}$-algebra $B$. In this case we will simply use the notation $A\otimes B$ to refer to the completion of $A\odot B$ in the necessarily unique $C^{*}$-norm on $A\odot B$.
By a trace $\tau$ on a $C^{*}$-algebra $B$ we will mean that $\tau$ is a nonzero tracial positive functional on $B$. In the event that $\tau$ is normalized, hence a state on $B$, we will call $\tau$ a tracial state. We make this distinction to ease the exposition, and in the case where we discuss matrix algebras, $M_{n}(\mathbb{C})$, the most natural trace to consider is the unnormalized one.
If $B$ is a $C^{*}$-algebra, we will denote by $B^{\textup{op}}=\{b^{\rm op}:b\in B\}$ the opposite $C^{*}$-algebra of $B$, with product given by $b^{\rm op}c^{\rm op}=(cb)^{\rm op}$ for all $b,c\in B$. The multiplication in $B$ induces a linear map $\mu\colon B\odot B^{\rm op}\to B$ defined by
$\mu(b_{1}\otimes b_{2}^{\rm op})=b_{1}b_{2}.$
In the event that $B$ admits a trace $\tau$, we can further compose $\mu$ with $\tau$ to get a linear map $\mu_{\tau}\colon B\odot B^{\rm op}\to\mathbb{C}$, which is determined by the following equation:
$\mu_{\tau}(b_{1}\otimes b_{2}^{\rm op})=\tau(b_{1}b_{2}).$
It is natural to ask for which $C^{*}$-norms on $B\odot B^{\rm op}$ the map $\mu_{\tau}$ is continuous. For the maximal norm we have the following result which follows by elementary methods, cf. [Oza04, Page 21] for an exposition.
Lemma 2.1.
Let $B$ be a $C^{*}$-algebra, and $\tau$ be a trace on $B$. Then $\mu_{\tau}$ is continuous in the maximal norm.
For the spatial norm, we have the following theorem due to Kirchberg.
Theorem 2.2 (Kirchberg).
Let $B$ be a unital $C^{*}$-algebra with a trace $\tau$. Then $\mu_{\tau}$ is continuous in the spatial norm if and only if $\tau$ is amenable.
This is only two of several equivalent statements proved by Kirchberg distributed over the papers [Kir93] and [Kir94]. A proof of these equivalences can be found in [BO08, Theorem 6.2.7]. Note that by amenability of $\tau$ we mean that the tracial state $\frac{1}{\tau(1_{B})}\tau$ is amenable in the following sense:
Definition 2.3 ([BO08]).
Let $A\subset B(H)$ be a concretely represented unital $C^{*}$-algebra. A state $\tau$ on $A$ is called an amenable trace if there exists a state $\varphi$ on $B(H)$ such that
  1. 1.
    $\varphi|_{A}=\tau$, and
  2. 2.
    $\varphi(uTu^{*})=\varphi(T)$ for every unitary $u\in A$ and $T\in B(H)$.
Remark 2.4.
The state $\varphi$ in Definition 2.3 is sometimes referred to as a hypertrace or an $A$-central state.
Remark 2.5.
Whenever $\mu_{\tau}$ is continuous with respect to some $C^{*}$-norm $\delta$, we will use the notation $\mu_{\tau}^{\delta}$ to denote the (unique) extension of $\mu_{\tau}$ to $B\otimes_{\delta}B^{\rm op}$. In the case where $B$ is nuclear, no extra decoration will be made to denote the extension of $\mu_{\tau}$ to $B\otimes B^{\rm op}$. In the case where $\delta=\|\cdot\|_{\rm min}$ or $\delta=\|\cdot\|_{\rm max}$, we will use the notation $\mu_{\tau}^{\rm min}$ or $\mu_{\tau}^{\rm max}$ respectively.
Note that in [BO08], $\mu_{\tau}$ is defined in a slightly different way, which makes it clear that $\mu_{\tau}$ is positive. We include a proof of this fact for the sake of completeness.
Lemma 2.6.
Let $B$ be a $C^{*}$-algebra with a trace $\tau$. Then the linear functional $\mu_{\tau}^{\rm max}$ is a positive functional on $B\otimes_{\rm max}B^{\rm op}$. Furthermore, if $B$ is unital and $\tau$ is amenable, then $\mu_{\tau}^{\rm min}$ is a positive functional on $B\otimes_{\rm min}B^{\rm op}$.
Proof.
Let $x=\sum_{j=1}^{n}b_{1,j}\otimes b_{2,j}^{\rm op}\in B\odot B^{\rm op}$. Then we compute that
$\mu_{\tau}(x^{*}x)$ $=\sum_{i,j=1}^{n}\mu_{\tau}(b_{1,j}^{*}b_{1,i}\otimes(b_{2,j}^{*})^{\rm op}b_{2,i}^{\rm op})=\sum_{i,j=1}^{n}\mu_{\tau}(b_{1,j}^{*}b_{1,i}\otimes(b_{2,i}b_{2,j}^{*})^{\rm op})$
$=\sum_{i,j=1}^{n}\tau(b_{1,j}^{*}b_{1,i}b_{2,i}b_{2,j}^{*})=\sum_{i,j=1}^{n}\tau(b_{2,j}^{*}b_{1,j}^{*}b_{1,i}b_{2,i})$
$=\sum_{i,j=1}^{n}\tau((b_{1,j}b_{2,j})^{*}b_{1,i}b_{2,i})=\tau\left(\sum_{j=1}^{n}(b_{1,j}b_{2,j})^{*}\sum_{i=1}^{n}(b_{1,i}b_{2,i})\right)$
$=\tau\left(\left(\sum_{j=1}^{n}b_{1,j}b_{2,j}\right)^{*}\left(\sum_{i=1}^{n}b_{1,i}b_{2,i}\right)\right).$
Hence, $\mu_{\tau}$ is positive on $B\odot B^{\rm op}$. Positivity on $B\otimes_{\rm max}B^{\rm op}$ follows from Lemma 2.1. In the case where $B$ is unital and $\tau$ is amenable, positivity of $\mu_{\tau}$ on $B\otimes_{\rm min}B^{\rm op}$ follows by Theorem 2.2, completing the proof. ∎
We are going to need the following result throughout this paper, cf. [BO08, Theorem 3.5.3] for a proof of this fact.
Lemma 2.7.
If $F\colon A\to B$ and $G\colon C\to D$ are completely positive maps between $C^{*}$-algebras, we have that the map
$F\odot G\colon A\odot C\to B\odot D$
extends to a completely positive map on both the minimal and maximal tensor products. Denoting (either) of the extensions by $F\otimes G$ we have have that
$\|F\otimes G\|=\|F\|\|G\|.$
We will often work with multiple tensor products and how these interact with opposite algebras. Here we establish some notation: Let $A$, $B$, and $C$ be $C^{*}$-algebras. Define $\Sigma\colon A\odot B\to B\odot A$ by
$\Sigma(a\otimes b)=b\otimes a.$
This is a *-isomorphism, hence $\Sigma$ is continuous in every $C^{*}$-norm on $A\odot B$. We will therefore not use any decoration to indicate the continuous extension of $\Sigma$, and we will furthermore not use any decoration to indicate the domain and codomain of $\Sigma$ either. If there are multiple tensor products present, we will use the subscript notation $\Sigma_{[ij]}$ to denote that we are interchanging the $i$-th and $j$-th tensor factor. For example, $\Sigma_{[23]}\colon A\odot B\odot C\to A\odot C\odot B$ is defined to be $\operatorname{id}_{A}\otimes\Sigma$.
Even though the map $B\to B^{\rm op}$, $b\mapsto b^{\rm op}$ does not necessarily define an anti-*-isomorphism between $B$ and $B^{\rm op}$, it is always a linear isomorphism. The assignment $b\mapsto b^{\rm op}$ is furthermore positive as
$b^{*}b\mapsto(b^{*}b)^{\rm op}=b^{\rm op}(b^{\rm op})^{*}.$
Hence the map $\Sigma^{\rm op}\colon A\odot B^{\rm op}\to B\odot A^{\rm op}$ defined by
$\Sigma^{\rm op}(a\otimes b^{\rm op})=b\otimes a^{\rm op}$
is a positive linear isomorphism, which furthermore is continuous in every $C^{*}$-norm on $A\odot B^{\rm op}$. We will again not decorate $\Sigma^{\rm op}$ to indicate that we are using the continuous extension, or to indicate the domain and codomain of $\Sigma^{\rm op}$. If there are multiple tensor factors we will again use the subscript notation $\Sigma_{[ij]}^{\rm op}$ to indicate that we are interchanging the $i$-th and $j$-th tensor factor and passing to the opposite algebra in the $i$-th and $j$-th tensor factor. For example, $\Sigma_{[23]}^{\rm op}\colon A\odot B\odot C^{\rm op}\to A\odot C\odot B^{\rm op}$ is defined to be $\operatorname{id}_{A}\otimes\Sigma^{\rm op}$.
We will use the notation $\operatorname{Hom}(A,B)$ to denote the vector space of all linear maps from $A$ to $B$, where we simply regard $A$ and $B$ as complex vector spaces. For the continuous dual of $A$ we will use the notation $A^{*}$, and if $F\colon A\to B$ is a bounded linear map we will denote the dual map by $F^{*}\colon B^{*}\to A^{*}$. The dual map is defined by setting
$F^{*}\psi=\psi\circ F,$
for any $\psi\in B^{*}$.
For a unital $C^{*}$-algebra $A$, we will denote the unit by $1_{A}$. In the special case where $A=M_{n}(\mathbb{C})$, we will for simplification use the notation $1_{n}$ instead of $1_{M_{n}(\mathbb{C})}$. Similarly, we will use the notation $\operatorname{id}_{n}$ for the identity map on $M_{n}(\mathbb{C})$.
2.2 Lipschitz seminorms and spectral triples
We remind the reader about basic notions from quantum metric geometry, in particular the definition of a Lipschitz seminorm. There are numerous definitions in the literature, e.g. [Rie04, Rie04a, Rie16, Kaa24], all of which are related and serve different purposes. In the sequel we will find it beneficial to use the definition from [KK25].
Definition 2.8.
Let $A$ be be a unital $C^{*}$-algebra. A Lipschitz seminorm on $A$ is a seminorm $L\colon A\to[0,\infty]$ satisfying
  1. 1.
    $L$ is densely defined in the sense that the domain ${\rm Dom}(L):=\{a\in A\mid L(a)<\infty\}$ is dense in $A$.
  2. 2.
    $L(a)=L(a^{*})$ for all $a\in A$.
  3. 3.
    $\mathbb{C}\cdot 1_{A}\subseteq\ker(L):=\{x\in A:L(x)=0\}$.
In general, by a seminorm on $A$, we will henceforth always mean a seminorm $L\colon A\to[0,\infty]$ that is densely defined in the sense above. Given such a seminorm $L$ on a unital $C^{*}$-algebra $A$, we may associate to it the Monge-Kantorovič extended metric $\operatorname{mk}_{L}\colon S(A)\times S(A)\to[0,\infty]$ given by
$\operatorname{mk}_{L}(\phi,\psi):=\sup\{|\phi(a)-\psi(a)|:a\in A,L(a)\leq 1\}$ (2.1)
When $L$ is a Lipschitz seminorm on $A$, there has been a lot of interest in deciding when $\operatorname{mk}_{L}$ metrizes the weak topology on $S(A)$, that is, when the induced topology from this extended metric coincides with the weak-topology. This is a central question when studying compact quantum metric spaces, but we will never explicitly need to consider this question in this article.
A rich source of examples of Lipschitz seminorms on unital $C^{*}$-algebras comes from spectral triples, of which we now remind the reader. Since we in Section 6.1 will closely follow the exposition of [Kaa24] on Kasparov products, we use the following definition of unital spectral triples, see [Con89, Con94, Kaa24].
Definition 2.9.
A unital spectral triple $(A,H,\partial)$ consists of a unital $C^{*}$-algebra $A$, a unital faithful representation $\pi\colon A\to B(H)$ on a separable Hilbert space $H$, and a self-adjoint (typically unbounded) operator $\partial\colon{\rm Dom}(\partial)\subseteq H\to H$ such that
  1. (a)
    there exists a dense *-subalgebra $\mathcal{A}\subseteq A$ such that for every $a\in\mathcal{A}$ the commutator
    $[\partial,\pi(a)]=\partial\pi(a)-\pi(a)\partial$
    is densely defined and extends to a bounded operator on $H$;
  2. (b)
    the resolvent $(i+\partial)^{-1}$ is a compact operator.
A unital spectral triple is even when the Hilbert space $H$ is equipped with a $\mathbb{Z}/2\mathbb{Z}$-grading operator $\gamma\colon H\to H$ such that $\pi(a)\colon H\to H$ is an even operator for all $a\in\mathcal{A}$ and $D\colon{\rm Dom}(D)\to H$ is odd. Otherwise we say the unital spectral triple is odd.
In the sequel we will suppress the *-homomorphism $\pi$ from the notation. To a spectral triple $(A,H,\partial)$ we may naturally associate a Lipschitz seminorm $L_{\partial}\colon A\to[0,\infty]$ through
$L_{\partial}(a)=\begin{cases}\|[\partial,a]\|&\text{if $a\in\mathcal{A}$}\\ \infty&\text{if $a\in A\setminus\mathcal{A}$},\end{cases}$ (2.2)
where we for notational convenience have identified $\partial,a]$ with its bounded extension. It is easy to verify that $L_{\partial}$ is a Lipschitz seminorm on $A$: The domain may be identified with $\mathcal{A}$, which is dense in $A$. The statement $\mathbb{C}\cdot 1_{A}\subseteq\ker(L_{\partial})$ follows from the unit commuting with $\partial$, and *-invariance follows from $\partial$ being self-adjoint.
Remark 2.10.
Constructing a Lipschitz seminorm from a spectral triple highlights an important subtlety: We set the domain ${\rm Dom}(L)$ to be equal to $\mathcal{A}$, even if it is possible that the quantity $\|[\partial,a]\|<\infty$ for some $a\in A\setminus\mathcal{A}$. Despite this, we still set $L_{\partial}(a)=\infty$ in this case. All this is to say that the domain of $L$ need not agree with the maximal domain for the derivation $a\mapsto[\partial,a]$.
We will in the sequel need to lift seminorms from $C^{*}$-algebras to their tensor products. Let $A$ and $B$ be unital $C^{*}$-algebras equipped with seminorms $L_{A}\colon A\to[0,\infty]$ and $L_{B}\colon B\to[0,\infty]$, respectively. Denote by ${\rm Dom}(L_{A})$ and ${\rm Dom}(L_{B})$ the domains of $L_{A}$ and $L_{B}$, respectively. Letting $A\otimes B$ denote either the minimal or maximal tensor product of $A$ and $B$, we define the following seminorms on $A\otimes B$
$\begin{split}(L_{A}\otimes 1)(z)&:=\sup\{L((\operatorname{id}_{A}\otimes\psi)(z))\colon\psi\in S(B)\}\\ (1\otimes L_{B})(z)&:=\sup\{L((\phi\otimes\operatorname{id}_{B})(z))\colon\phi\in S(A)\}\\ L_{A\otimes B}&:=L_{A}\otimes 1+1\otimes L_{B}\end{split}$ (2.3)
where the domain for all three of these seminorms is set to be ${\rm Dom}(L_{A})\odot{\rm Dom}(L_{B})$ which is dense in $A\otimes B$. Assume now that $L_{A}$ and $L_{B}$ are Lipschitz seminorms. Using the fact that $L_{A}$ and $L_{B}$ are both *-invariant, it is not difficult to verify that the three seminorms just defined are *-invariant. Moreover, we easily verify that $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$ is contained in the kernel of all three seminorms. Thus they are all Lipschitz seminorms. Note however that $\ker(L_{A}\otimes 1)$ is generally larger than $\mathbb{C}\cdot(1_{A}\otimes 1_{B})$, as it at the very least contains $\mathbb{C}\cdot(1_{A}\otimes{\rm Dom}(L_{B}))$. The analogous statement holds for $1\otimes L_{B}$. Moreover, $L_{A\otimes B}(x)=L_{B\otimes A}(\Sigma(x))$ for all $x\in A\otimes B$.
Lastly, we will at times need to pass from a $C^{*}$-algebra $A$ to its opposite $C^{*}$-algebra $A^{\rm op}$. When $A$ comes equipped with a seminorm $L_{A}$ we induce a seminorm $L_{A^{\rm op}}$ on $A^{\rm op}$ defined by
$L_{A^{\rm op}}(a^{\rm op}):=L_{A}(a)$ (2.4)
from which we also note that $a^{\rm op}\in{\rm Dom}(L_{A^{\rm op}})\subseteq A^{\rm op}$ if and only if $a\in{\rm Dom}(L_{A})\subseteq A$. Clearly, $L_{A^{\rm op}}$ is a Lipschitz seminorm if and only if $L_{A}$ is a Lipschitz seminorm.
3 Discussions on inducing metrics on unital completely positive maps
Throughout this section, $A$ and $B$ will be unital $C^{*}$-algebras, and we will be considering extended metrics on $\operatorname{UCP}(A,B)$, the set of unital completely positive maps from $A$ to $B$. In [GLN05] and [BV25], several desirable properties of such metrics are discussed, as well as some examples of metrics in the case where $A$ and $B$ are matrix algebras. Note that in [GLN05, BV25] they discuss quantum channels between $A$ and $B$ instead of unital completely positive maps. These classes of maps are dual to each other, and for the sake of clarity, we will simply translate the discussions in [GLN05] and [BV25] to how it would fit into our current setting without further mentions.
The first property an extended metric on $\operatorname{UCP}(A,B)$ should satisfy is stability. In both [GLN05] and [BV25], a metric $d$ on $\operatorname{UCP}(A,B)$ is said to be stable if
$d(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d(F,G)$ (3.1)
holds for every $n\in\mathbb{N}$ and $F,G\in\operatorname{UCP}(A,B)$. From a quantum information theoretic perspective, this property is supposed to reflect the fact that an unrelated ancillary system should not affect the distance between the unital completely positive maps. Note however that Equation 3.1 does not make sense on the nose, because as it currently reads $d$ is a metric not only on the unital completely positive maps between $A$ and $B$, but also simultaneously on their amplifications. In [GLN05] and [BV25] it is expected that an analogue of $d$ exists on $\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ for every $n\in\mathbb{N}$. For our purposes this need not be the case, and we therefore define stability as a property of a sequence of metrics.
Definition 3.1 (Stability).
Let $A$ and $B$ be unital $C^{*}$-algebras, and $\{d_{n}\}_{n\in\mathbb{N}}$ be sequence of extended metrics where each $d_{n}$ is an extended metric on
$\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B).$
We say that the sequence $\{d_{n}\}_{n\in\mathbb{N}}$ is stable if
$d_{n}(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)=d_{1}(F,G)$ (3.2)
for every $n\in\mathbb{N}$ and $F,G\in\operatorname{UCP}(A,B)$.
In Definition 4.13 we will introduce the set $\operatorname{TC}_{\tau}(A,B)$ of trace channels from $A$ to $B$, where $A$ is a unital $C^{*}$-algebra, and $B$ is a $C^{*}$-algebra with a faithful trace $\tau$. These form a particular subset of the completely positive maps from $A$ to $B$. We will then instead analogously define stability for a sequence $\{d_{n}\}_{n\in\mathbb{N}}$, where $d_{n}$ is an extended metric on
$\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B).$
The second property is chaining, which aims to capture the idea that if one divides a process into smaller steps, the total error should be smaller than the sum of the step-wise errors. Mathematically, it is defined as follows.
Definition 3.2 (Chaining).
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras. We say that the three metrics $d_{1}$, $d_{2}$, and $d_{3}$ defined on $\operatorname{UCP}(A,B)$, $\operatorname{UCP}(B,C)$, and $\operatorname{UCP}(A,C)$ respectively, satisfy chaining if given $F_{1},F_{2}\in\operatorname{UCP}(A,B)$, and $G_{1},G_{2}\in\operatorname{UCP}(B,C)$, the following inequality is satisfied:
$d_{3}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq d_{2}(G_{1},G_{2})+d_{1}(F_{1},F_{2}).$ (3.3)
In the special case where $A=B=C$ and $d_{1}=d_{2}=d_{3}=d$, we say that $d$ satisfies chaining.
Once again we will want to work with trace channels. It will however turn out that the composition of two trace channels is not necessarily a trace channel, hence we will in the sequel need to be careful when generalizing the definition of chaining to include trace channels as well. We will get back to this in Section 5.2.
In [GLN05] and [BV25] they discuss different ways of inducing metrics on $\operatorname{UCP}(A,B)$ given a metric on $S(A)$, or a metric on $S(A\otimes B^{\rm op})$ (recall that they only consider matrix algebras, hence there is a unique tensor product of $C^{*}$-algebras in this case). However, instead of starting with a metric on any of these state spaces, we start with noncommutative geometric data in the form of a seminorm. One way of inducing an extended metric on $\operatorname{UCP}(A,B)$ discussed in [GLN05, BV25] is as follows.
Definition 3.3.
Let $A$ and $B$ be unital $C^{*}$-algebras, and $L\colon A\to[0,\infty]$ be any seminorm on $A$. Define the extended metric $D_{L}\colon\operatorname{UCP}(A,B)\times\operatorname{UCP}(A,B)\to[0,\infty]$ by setting
$D_{L}(F,G):=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(F^{*}\psi,G^{*}\psi).$ (3.4)
Note that if we have a sequence $\{L_{n}\}_{n\in\mathbb{N}}$ of seminorms $L_{n}\colon M_{n}(\mathbb{C})\otimes A\to[0,\infty]$, it is generally not true that $\{D_{L_{n}}\}_{n\in\mathbb{N}}$ will be a stable sequence of extended metrics. Even if $L_{n}=\|\cdot\|$, the operator norm on $M_{n}(\mathbb{C})\otimes A$, the sequence $\{D_{L_{n}}\}_{n\in\mathbb{N}}$ will not be stable. A way to fix this problem is as follows: Let
$D_{n}^{\rm stab}(F,G):=\sup\limits_{m\in\mathbb{N}}D_{L_{mn}}(\operatorname{id}_{mn}\otimes F,\operatorname{id}_{mn}\otimes G),$ (3.5)
where $L_{mn}$ is still the operator norm on $M_{mn}(\mathbb{C})\otimes A$. Then $\{D_{n}^{\rm stab}\}_{n\in\mathbb{N}}$ is stable. This is analogous to how the diamond norm, and the completely bounded norm are defined. Even though this sequence of metrics satisfies stability, it has the drawback that explicit calculations depend on calculating distances over all the matrix amplifications. Moreover, the construction only takes into consideration the geometric structure of $A$, not the target $C^{*}$-algebra $B$, therefore ignoring potentially important data present in the situation.
Another way of inducing a metric on $\operatorname{UCP}(A,B)$ discussed in [GLN05] and [BV25] involves embedding $\operatorname{UCP}(A,B)$ into the state space of a higher-dimensional $C^{*}$-algebra using the Choi-Jamiolkowski isomorphism. In Section 4 we consider a $C^{*}$-algebra $B$ with a faithful trace $\tau$, and we define the map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$ which gives an infinite-dimensional analogue of the map appearing in the Choi-Jamiolkowski isomorphism. We show that this gives an embedding $\omega_{\tau}^{\rm max}$ of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm max}B^{\rm op})$. Furthermore, if $B$ is unital and $\tau$ is an amenable trace, then this gives an embedding $\omega_{\tau}^{\rm min}$ of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm min}B^{\rm op})$, cf. Theorem 4.15. With this we can make the following definitions.
Definition 3.4.
Let $A$ be a unital $C^{*}$-algebra and $B$ be a $C^{*}$-algebra with a trace $\tau$. Let $L\colon A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ be a seminorm on $A\otimes_{\rm max}B^{\rm op}$. Define the extended metric $\Delta_{\tau,L}^{\rm max}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
$\Delta_{\tau,L}^{\rm max}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm max}(F),\omega_{\tau}^{\rm max}(G)).$ (3.6)
In the case where $B$ is unital, $\tau$ is amenable, and $L\colon A\otimes_{\rm min}B^{\rm op}\to[0,\infty]$ is a seminorm on $A\otimes_{\rm min}B^{\rm op}$, we may define $\Delta_{\tau,L}^{\rm min}\colon\operatorname{TC}_{\tau}(A,B)\times\operatorname{TC}_{\tau}(A,B)\to[0,\infty]$ by setting
$\Delta_{\tau,L}^{\rm min}(F,G):=\operatorname{mk}_{L}(\omega_{\tau}^{\rm min}(F),\omega_{\tau}^{\rm min}(G)).$ (3.7)
In the rest of this paper we have chosen to focus on the metrics $\Delta_{\tau,L}^{\rm max}$ and $\Delta_{\tau,L}^{\rm min}$. The reason for this is that we are able to show, that for a sequence of seminorms $\{L_{n}\}_{n\in\mathbb{N}}$ (defined on the appropriate spaces) satisfying certain mild assumptions, the induced sequences $\{\Delta_{\tau,L_{n}}^{\rm max}\}_{n\in\mathbb{N}}$ and $\{\Delta_{\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$ will be stable, cf. Theorem 5.3. In Section 6 we show that the assumptions needed for stability are satisfied for a large class of $C^{*}$-algebras when the seminorms are induced from spectral triples in an appropriate manner. Moreover, we also show that chaining for both $\Delta_{\tau,L}^{\rm max}$ and $\Delta_{\tau,L}^{\rm min}$ also follow for subsets of the set of trace channels under mild assumptions on $L$, cf. Theorems 5.10 and 5.12. Note that in a lot of circumstances it might be too strict to require that Equation 3.3 holds for all unital completely positive maps, or all trace channels. Indeed, even in the matrix algebra case, [GLN05] find it necessary to put restrictions on at least one of the maps involved in their analogue of Equation 3.3, see [GLN05, pg. 5]. We manage to prove that chaining is satisfied for a large class of unital completely positive maps between twisted group $C^{*}$-algebras when the seminorms are induced from spectral triples arising from proper length functions on the underlying group, cf. Theorem 6.9.
4 Tracial $C^{*}$-algebras and completely positive maps
In this section we consider a general $C^{*}$-algebra $A$ and a $C^{*}$-algebra $B$ with a trace $\tau$. We will define a map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$, which we will use to characterize the linear maps $F\colon A\to B$ satisfying that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm max}B^{\rm op}$. We will call these maps trace channels. In the case where $B$ is unital and $\tau$ is amenable, Kirchberg’s Theorem, Theorem 2.2, will give that $\omega_{\tau}(F)$ extends to a state on $A\otimes_{\rm min}B^{\rm op}$ as well.
4.1 The map $\omega_{\tau}$ and trace channels
Definition 4.1.
Let $A$ and $B$ be $C^{*}$-algebras, and suppose that $\tau$ is a trace on $B$. For any linear map $F\colon A\to B$ we define the linear functional $\omega_{\tau}(F)\colon A\odot B^{\rm op}\to\mathbb{C}$ by setting
$\omega_{\tau}(F)(a\otimes b^{\rm op})=\tau(F(a)b)$ (4.1)
for any $a\in A$ and $b\in B$. By the universal property of the algebraic tensor product, this uniquely defines $\omega_{\tau}(F)$ as a linear functional on $A\odot B^{\rm op}$. It is clear that the assignment $F\mapsto\omega_{\tau}(F)$ is linear, hence we get a linear map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$.
Remark 4.2.
Note that $\omega_{\tau}(F)$ is given by the following composition
${A\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F\odot\operatorname{id}_{B^{\rm op}}}$${B\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\tau}}$${\mathbb{C}.}$
We will often make use of this fact.
We are furthermore interested in how $\omega_{\tau}$ interacts with different $C^{*}$-norms on $A\odot B^{\rm op}$. In particular, if $\delta$ is a $C^{*}$-norm on $A\odot B^{\rm op}$, we want to now for which linear maps $F\colon A\to B$ do we have that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. To that end, we give the following definition.
Definition 4.3.
Let $A$ and $B$ be $C^{*}$-algebras, and suppose that $\tau$ is a trace on $B$. For a $C^{*}$-norm $\delta$ on $A\odot B^{\rm op}$ define the set $\mathcal{L}_{\tau}^{\delta}(A,B)$ to be the set of linear maps $F\colon A\to B$ such that $\omega_{\tau}(F)$ is continuous with respect to $\delta$. In this case we use the notation $\omega_{\tau}^{\delta}(F)$ to denote the (unique) extension of $\omega_{\tau}(F)$ to $A\otimes_{\delta}B^{\rm op}$. If either $A$ or $B$ is nuclear, no extra decoration will be made to denote the extension of $\omega_{\tau}(F)$ to $A\otimes B^{\rm op}$.
In the special cases where $\delta=\|\cdot\|_{\rm min}$ and $\delta=\|\cdot\|_{\rm max}$ we will simply use the notation $\mathcal{L}_{\tau}^{\rm min}(A,B)$ and $\mathcal{L}_{\tau}^{\rm max}(A,B)$ respectively.
Remark 4.4.
In this definition we are considering the possibility that $\omega_{\tau}(F)$ may be continuous in some $C^{*}$-norm on $A\odot B$ without assuming anything about the continuity of $\mu_{\tau}$. If $\mu_{\tau}$ is continuous (as it always is with respect to the maximal norm), then continuity of $\omega_{\tau}(F)$ follows by continuity of $F\odot\operatorname{id}_{B^{\rm op}}$.
Proposition 4.5.
Let $A$ and $B$ be $C^{*}$-algebras. Suppose that $\tau$ is a trace on $B$ and that $\delta$ is a $C^{*}$-norm on $A\odot B^{\rm op}$. Then,
  1. 1.
    $\mathcal{L}_{\tau}^{\delta}(A,B)$ is a subspace of $\operatorname{Hom}(A,B)$,
  2. 2.
    the assignment $F\mapsto\omega_{\tau}^{\delta}(F)$ defines a linear map $\omega_{\tau}^{\delta}\colon\mathcal{L}_{\tau}^{\delta}(A,B)\to(A\otimes_{\delta}B^{\rm op})^{*}$, and
  3. 3.
    if $\tau$ is faithful, then the map $\omega_{\tau}^{\delta}$ is a linear embedding.
Proof.
Let $F\in\mathcal{L}_{\tau}^{\delta}(A,B)$. That $\mathcal{L}_{\tau}^{\delta}(A,B)$ is a vector space follows by the fact that $\omega_{\tau}$ is linear.
That $\omega_{\tau}^{\delta}$ is linear follows by linearity of $\omega_{\tau}$ and the fact that $\omega_{\tau}^{\delta}(F)$ is the unique extension of $\omega_{\tau}(F)$ to $A\otimes_{\delta}B^{\rm op}$.
Suppose now that $\tau$ is faithful. We want to show that $\omega_{\tau}^{\delta}(F)=0$ if and only if $F=0$. Since $\omega_{\tau}^{\delta}(F)$ is the unique extension of $\omega_{\tau}(F)$ it is sufficient to show that $\omega_{\tau}(F)=0$ if and only if $F=0$. Suppose that $\omega_{\tau}(F)=0$. Then for any $a\in A$ and $b\in B$ we have that $\tau(F(a)b)=0$. Since $\tau$ is faithful this means that $F(a)=0$ for every $a\in A$, hence $F=0$. ∎
We will show that the completely positive maps from $A$ to $B$ are precisely the maps $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ for which $\omega_{\tau}^{\rm max}(F)$ is positive. First we show that if $F\in\operatorname{CP}(A,B)$, then $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$.
Proposition 4.6.
Let $A$ and $B$ be $C^{*}$-algebras, and suppose that $\tau$ is a trace on $B$. Then $\operatorname{CP}(A,B)\subset\mathcal{L}_{\tau}^{\rm max}(A,B)$.
Furthermore, if $B$ is unital and $\tau$ is amenable, then $\operatorname{CP}(A,B)\subset\mathcal{L}_{\tau}^{\rm min}(A,B)$.
Proof.
In the case that $F$ is completely positive, we have by Lemma 2.7 and Lemma 2.1 that $\omega_{\tau}(F)$ is continuous in the maximal norm. This proves the first claim.
In the case that $B$ is unital and $\tau$ is an amenable trace, we get by Lemma 2.7 and Theorem 2.2 that $\omega_{\tau}(F)$ is continuous in the spatial norm. ∎
Remark 4.7.
If $B$ is quasidiagonal, then $B$ has at least one amenable tracial state $\tau$, cf. [BO08, Proposition 7.1.16], so $\omega_{\tau}(F)$ is continuous in the spatial norm on $A\odot B^{\rm op}$ for every $F$ in $\operatorname{CP}(A,B)$ in this case. For examples of unital, separable, non-nuclear, quasidiagonal $C^{*}$-algebras, see [Pis20, Remark 7.3].
Remark 4.8.
Whenever $F\in{\rm Dec(A,B)}:={\rm Span}(\operatorname{CP}(A,B))$ we have that $\omega_{\tau}(F)$ is continuous in the maximal norm on $A\odot B^{\rm op}$. This follows in the same way as in the proof of Proposition 4.6, now using [Pis20a, Proposition 6.11], which implies that
$F\odot\textup{id}_{B^{\textup{op}}}\colon A\odot B^{\textup{op}}\to B\odot B^{\textup{op}}$
extends to a map
$F\otimes\textup{id}_{B^{\textup{op}}}\colon A\otimes_{\rm max}B^{\textup{op}}\to B\otimes_{\rm max}B^{\textup{op}}$
for every $F\in{\rm Dec(A,B)}$. On the other hand, if we assume that $\tau$ is amenable, then we may use [Pis20a, Proposition 1.11] to deduce that $\omega_{\tau}(F)$ is continuous in the spatial norm on $A\odot B^{\rm op}$ for every $F\in\operatorname{CB}(A,B)$, where $\operatorname{CB}(A,B)$ denotes the set of all completely bounded maps from $A$ to $B$.
Remark 4.9.
A sufficient condition for the continuity of $\omega_{\tau}(F)$ in the spatial norm on $A\odot B^{\rm op}$ depending only on $F$ is as follows. Assume $F$ is the pointwise limit of a net of finite rank completely positive maps from $A$ to $B$. Then it follows from [Pis20a, Theorem 7.11] that the map
$F\odot\textup{id}_{B^{\textup{op}}}\colon A\odot B^{\textup{op}}\to B\odot B^{\textup{op}}$
extends to a continuous map
$F\otimes\textup{id}_{B^{\textup{op}}}\colon A\otimes_{\rm min}B^{\textup{op}}\to B\otimes_{\rm max}B^{\textup{op}}.$
As we know from Lemma 2.1, $\mu_{\tau}$ is continuous on $B\odot B^{\textup{op}}$ with respect to the maximal norm, so we get that $\omega_{\tau}(F)$ is continuous with respect to the spatial norm on $A\odot B^{\rm op}$ as it is given by the composition of the following continuous maps:
${A\otimes_{\rm min}B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F\otimes\operatorname{id}_{B^{\rm op}}}$${B\otimes_{\rm max}B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\tau}^{\rm max}}\textstyle{\mathbb{C}}$
Hence, $F\in\mathcal{L}_{\tau}^{\rm min}(A,B)$ in this case.
To prove the desired characterization of completely positive maps from $A$ to $B$, we need to understand how the map $\omega_{\tau}$ interacts with tensor products of linear maps.
Lemma 4.10.
Let $A$ and $C$ be $C^{*}$-algebras, and $B$ and $D$ be $C^{*}$-algebras with traces $\tau_{B}$ and $\tau_{D}$ respectively. Then for any $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$ we have that
$(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G).$ (4.2)
Furthermore, if $B$ and $D$ are unital, and $\tau_{B}$ and $\tau_{D}$ are amenable, then
$(\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G).$ (4.3)
Proof.
By Lemma 2.7 and Proposition 4.6 we have that the all the maps in the following diagram are well-defined.
${\operatorname{CP}(A,B)\times\operatorname{CP}(C,D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$$\scriptstyle{\omega_{\tau_{B}}^{\rm max}\times\omega_{\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}B^{\rm op})^{*}\times(C\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\otimes}$${\operatorname{CP}_{\tau_{B}\otimes\tau_{D}}(A\otimes_{\rm max}C,B\otimes_{\rm max}D)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}}$${(A\otimes_{\rm max}C\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}D^{\rm op})^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\Sigma_{[23]})^{*}}$${(A\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}C\otimes_{\rm max}D^{\rm op})^{*}}$ (4.4)
We have that this diagram commutes if and only if Equation 4.2 hold for all $F\in\operatorname{CP}(A,B)$ and $G\in\operatorname{CP}(C,D)$. Hence, we simply want to show that this diagram commutes. For $a\in A$, $b\in B$, $c\in C$, and $d\in D$ we have that
$((\Sigma_{[23]})^{*}\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G))(a\otimes b^{\rm op}\otimes c\otimes d^{\rm op})$ $=\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)(a\otimes c\otimes b^{\rm op}\otimes d^{\rm op})$
$=(\tau_{B}\otimes\tau_{D})((F(a)\otimes G(c))(b\otimes d))$
$=(\tau_{B}\otimes\tau_{D})(F(a)b\otimes G(c)d)$
$=\tau_{B}(F(a)b)\tau_{D}(G(c)d)$
$=\omega_{\tau_{B}}^{\rm max}(F)(a\otimes b^{\rm op})\omega_{\tau_{D}}^{\rm max}(G)(c\otimes d^{\rm op})$
$=(\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G))(a\otimes b^{\rm op}\otimes c\otimes d^{\rm op}).$
By linearity and continuity of the relevant maps, and the fact that $A\odot B^{\rm op}\odot C\odot D^{\rm op}$ is dense in $A\otimes_{\rm max}B^{\rm op}\otimes_{\rm max}C\otimes_{\rm max}D^{\rm op}$, we get that the above diagram commutes.
Now suppose that $B$ and $D$ are unital, and that $\tau_{B}$ and $\tau_{D}$ are amenable. By [BO08, Proposition 6.3.5] we get that $\tau_{B}\otimes\tau_{D}$ is an amenable trace on $B\otimes_{\rm min}D$. Hence, everything appearing in the diagram presented in Equation 4.4 is still well-defined with min in place of max by Lemma 2.7 and Proposition 4.6. That Equation 4.3 holds then follows by the same computation we did earlier in this proof. ∎
Remark 4.11.
Since the *-isomorphism $\Sigma_{[23]}$ is idempotent, we have that Equation 4.2 and Equation 4.3 give rise to analogous equations
$\omega_{\tau_{B}\otimes\tau_{D}}^{\rm max}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm max}(F)\otimes\omega_{\tau_{D}}^{\rm max}(G)),$ (4.5)
and
$\omega_{\tau_{B}\otimes\tau_{D}}^{\rm min}(F\otimes G)=(\Sigma_{[23]})^{*}(\omega_{\tau_{B}}^{\rm min}(F)\otimes\omega_{\tau_{D}}^{\rm min}(G)).$ (4.6)
We will also have use for these equations.
Theorem 4.12.
Let $A$ and $B$ be a $C^{*}$-algebra, and let $\tau$ be a faithful trace on $B$. For any $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ we have that $F$ is completely positive if and only if $\omega_{\tau}^{\rm max}(F)$ is positive.
Furthermore, if $B$ is unital, $\tau$ is amenable, and $F\in\mathcal{L}_{\tau}^{\rm min}(A,B)$, then $F$ is completely positive if and only if $\omega_{\tau}^{\rm min}(F)$ is positive.
Proof.
Suppose first that $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$ is completely positive. Since $\omega_{\tau}^{\rm max}(F)$ and $\mu_{\tau}^{\rm max}\circ(F\otimes\operatorname{id}_{B^{\rm op}})$ agree on the dense subspace $A\odot B^{\rm op}$ we have that $\omega_{\tau}^{\rm max}(F)$ is given by the following composition:
${A\otimes_{\rm max}B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F\otimes\operatorname{id}_{B^{\rm op}}}$${B\otimes_{\rm max}B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{\tau}^{\rm max}}$${\mathbb{C}.}$
The first of these maps is completely positive by Lemma 2.7, and the second map is positive by Lemma 2.6. Hence, $\omega_{\tau}^{\rm max}(F)$ is a composition of positive maps, therefore $\omega_{\tau}^{\rm max}(F)$ is positive.
For the converse, assume that $\omega_{\tau}^{\rm max}(F)$ is positive. Let $n\in\mathbb{N}$ and fix $Y\in M_{n}(\mathbb{C})\otimes_{\rm max}B$ to be positive. It follows that the map $\operatorname{Tr}_{n}\otimes\tau$ is a faithful trace on $M_{n}(\mathbb{C})\otimes_{\rm max}B$, cf. [Avi82, p. 434], as $M_{n}(\mathbb{C})\otimes_{\rm max}B\simeq M_{n}(\mathbb{C})\otimes_{\rm min}B$. Hence, to see that $F$ is $n$-positive it is enough to check that
$(\operatorname{Tr}_{n}\otimes\tau)((\operatorname{id}_{n}\otimes F)(X)Y)$
is positive for every positive element $X\in M_{n}(\mathbb{C})\otimes_{\rm max}A$ and every positive element $Y\in M_{n}(\mathbb{C})\otimes_{\rm max}B$. We compute that
$(\operatorname{Tr}_{n}\otimes\tau)((\operatorname{id}_{n}\otimes F)(X)Y)$ $=(\omega_{\operatorname{Tr}_{n}\otimes\tau}(\operatorname{id}_{n}\otimes F))(X\otimes Y^{\rm op})$ by Equation 4.1
$=(\omega_{\operatorname{Tr}_{n}\otimes\tau}^{\rm max}(\operatorname{id}_{n}\otimes F))(X\otimes Y^{\rm op})$
$=((\Sigma_{[23]})^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)))(X\otimes Y^{\rm op})$ by Equation 4.5
$=((\Sigma_{[23]})^{*}(\mu_{\operatorname{Tr}_{n}}\otimes\omega_{\tau}^{\rm max}(F)))(X\otimes Y^{\rm op})$
$=((\mu_{\operatorname{Tr}_{n}}\otimes\omega_{\tau}^{\rm max}(F))\circ\Sigma_{[23]})(X\otimes Y^{\rm op}).$
Note that $X\otimes Y^{\rm op}$ is positive since both $X$ and $Y$ are assumed to be positive. Recall that $\Sigma_{[23]}$ is a *-isomorphism, which in particular means that $\Sigma_{[23]}$ is positive. By Lemma 2.6 we have that $\mu_{\operatorname{Tr}_{n}}$ is positive, and we have assumed that $\omega_{\tau}^{\rm max}(F)$ is positive. Thus, by Lemma 2.7 we have that $\mu_{\operatorname{Tr}_{n}}\otimes\omega_{\tau}^{\rm max}(F)$ is positive. Since the composition of positive maps is positive, we have that
$(\operatorname{Tr}_{n}\otimes\tau)((\operatorname{id}_{n}\otimes F)(X)Y)$
is positive. This shows that $F$ is $n$-positive. But since $n$ was arbitrary, $F$ is necessarily completely positive.
In the case where $B$ is unital and $\tau$ is amenable, the argument given above works mutatis mutandis to show that any $F\in\mathcal{L}_{\tau}^{\rm min}(A,B)$ is completely positive if and only if $\omega_{\tau}^{\rm min}(F)$ is positive. ∎
By Theorem 4.12 we have that the linear maps from $A$ to $B$ that will correspond to states on $A\otimes_{\rm max}B^{\rm op}$ under the map $\omega_{\tau}^{\rm max}$ must necessarily be completely positive. To get a precise characterization in Theorem 4.15 we make the following definition.
Definition 4.13.
Let $A$ be a unital $C^{*}$-algebra and $B$ be a $C^{*}$-algebra with a trace $\tau$. Define the set
$\operatorname{TC}_{\tau}(A,B)=\{F\in\operatorname{CP}(A,B):\tau(F(1_{A}))=1\}.$
We will call elements of $\operatorname{TC}_{\tau}(A,B)$ trace channels (with respect to $\tau$).
We note that if $B$ is unital and $\tau_{B}$ is a tracial state, then $\operatorname{UCP}(A,B)$ is contained in $\operatorname{TC}_{\tau}(A,B)$.
Lemma 4.14.
Let $A$ be a unital $C^{*}$-algebra and $B$ be a $C^{*}$-algebra with a trace $\tau$. Then $\operatorname{TC}_{\tau}(A,B)$ is a closed convex subset of $\operatorname{CP}(A,B)$.
Proof.
Note that the condition $\tau(F(1_{A}))=1$ is a closed condition, hence $\operatorname{TC}_{\tau}(A,B)$ is a closed subset of $\operatorname{CP}(A,B)$. If $t\in(0,1)$ and $F,G\in\operatorname{TC}_{\tau}(A,B)$, then tF+(1t)GtF+(1-t)G is completely positive, since $\operatorname{CP}(A,B)$ is a convex set. By linearity of $\tau$ we also get that
$\tau((tF+(1-t)G)(1_{A}))=t\tau(F(1_{A}))+(1-t)\tau(G(1_{A}))=t+(1-t)=1.$
Hence, $tF+(1-t)G\in\operatorname{TC}_{\tau}(A,B)$, which is what we wanted to show. ∎
We can now state the main result of this section.
Theorem 4.15.
Let $A$ be a unital $C^{*}$-algebra, $B$ be a $C^{*}$-algebra with a faithful trace $\tau$, and $F\in\mathcal{L}_{\tau}^{\rm max}(A,B)$. Then the following are equivalent,
  1. 1.
    $\omega_{\tau}^{\rm max}(F)\in S(A\otimes_{\rm max}B^{\rm op})$, and
  2. 2.
    $F\in\operatorname{TC}_{\tau}(A,B)$.
In particular, $\omega_{\tau}^{\rm max}$ restricts to an affine embedding of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm max}B^{\rm op})$.
Furthermore, if $B$ is unital, $\tau$ is amenable, and $F\in\mathcal{L}_{\tau}^{\rm min}(A,B)$, then the following are equivalent,
  1. 1.
    $\omega_{\tau}^{\rm min}(F)\in S(A\otimes_{\rm min}B^{\rm op})$, and
  2. 2.
    $F\in\operatorname{TC}_{\tau}(A,B)$.
In this case, $\omega_{\tau}^{\rm min}$ restricts to an affine embedding of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm min}B^{\rm op})$.
Proof.
By Theorem 4.12 we have that $\omega_{\tau}^{\rm max}(F)$ is positive if and only if $F$ is completely positive. In this case, we need only show that $\omega_{\tau}^{\rm max}(F)$ is a state if and only if $\tau(F(1_{A}))=1$. Let $(u_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit for $B$. Then $(u_{\lambda}^{\rm op})_{\lambda\in\Lambda}$ is an approximate unit for $B^{\rm op}$. Note that
$\|\omega_{\tau}^{\rm max}(F)\|=\lim\limits_{\lambda}(\omega_{\tau}^{\rm max}(F)(1_{A}\otimes u_{\lambda}^{\rm op}))=\lim\limits_{\lambda}(\tau(F(1_{A})u_{\lambda}))=\tau(F(1_{A})),$
hence if $F$ is completely positive, we have that $\omega_{\tau}^{\rm max}(F)$ is a state if and only if $\tau(F(1_{A}))=1$. That $\omega_{\tau}^{\rm max}$ restricts to an affine embedding of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm max}B^{\rm op})$ now follows by Lemma 4.14.
The proof for the case when $B$ is unital and $\tau$ is amenable follows by exactly the same argument. ∎
Definition 4.16.
Let $A$ and $B$ be $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$ respectively. We say that a linear map $F\colon A\to B$ is $(\tau_{A},\tau_{B})$-adjointable if there exists a linear map $F^{\sharp}\colon B\to A$ such that
$\tau_{B}(F(a)b)=\tau_{A}(aF^{\sharp}(b))$ (4.7)
for every $a\in A$ and $b\in B$.
Proposition 4.17.
Let $A$ and $B$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$ respectively. If $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ is $(\tau_{A},\tau_{B})$-adjointable, then $F^{\sharp}\in\operatorname{TC}_{\tau_{A}}(B,A)$, and we furthermore have that
$\omega_{\tau_{A}}(F^{\sharp})=\omega_{\tau_{B}}(F)\circ\Sigma^{\rm op}.$ (4.8)
Proof.
First note that
$\tau_{A}(F^{\sharp}(1_{B}))=\tau_{A}(1_{A}F^{\sharp}(1_{B}))=\tau_{B}(F(1_{A})1_{B})=\tau_{B}(F(1_{A}))=1.$
Hence, it remains to show that $F^{\sharp}$ is completely positive. For $a\in A$ and $b\in B$ we have that
$\omega_{\tau_{A}}(F^{\sharp})(b\otimes a^{\rm op})$ $=\tau_{A}(F^{\sharp}(b)a)$
$=\tau_{A}(aF^{\sharp}(b))$
$=\tau_{B}(F(a)b)$
$=\omega_{\tau_{B}}(F)(a\otimes b^{\rm op})$
$=(\omega_{\tau_{B}}(F)\circ\Sigma^{\rm op})(b\otimes a^{\rm op}).$
Hence, the following diagram commutes:
${B\odot A^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\tau_{A}}(F^{\sharp})}$$\scriptstyle{\Sigma^{\rm op}}$${A\odot B^{\rm op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\tau_{B}}(F)}\textstyle{\mathbb{C}}$
By Proposition 4.6 we have that $F\in\mathcal{L}_{\tau_{B}}^{\rm max}(A,B)$. Hence, both $\Sigma^{\rm op}$ and $\omega_{\tau_{B}}(F)$ are continuous in the maximal norm, thus $\omega_{\tau_{A}}(F^{\sharp})$ necessarily extends to a continuous map on the maximal tensor product. This shows that $F^{\sharp}\in\mathcal{L}_{\tau_{A}}^{\rm max}(B,A)$.
By Theorem 4.12 we furthermore have that $\omega_{\tau_{B}}^{\rm max}(F)$ is positive. Recall that $\Sigma^{\rm op}$ is also positive, hence $\omega_{\tau_{A}}^{\rm max}(F^{\sharp})$ is positive. By Theorem 4.12 we get that $F^{\sharp}$ is completely positive. This shows that $F^{\sharp}\in\operatorname{TC}_{\tau_{A}}(B,A)$. ∎
4.2 Relation to the Choi-Jamiolkowski isomorphisms
Theorem 4.12 can be used to recover some characterizations of completely positive maps that are known for matrix algebras and finite-dimensional $C^{*}$-algebras. We discuss some of these relationships in this subsection.
In the case that a $C^{*}$-algebra $B$ admits a faithful trace $\tau$, we get a linear embedding of $B$ into its dual space $B^{*}$ by the assignment $b\mapsto\varphi_{b}:=\tau(b\,\cdot\,)$. When $B$ is finite-dimensional, this is a linear isomorphism, and the elements of
$D_{\tau}(B):=\{b\in B:b\text{ is positive and }\tau(b)=1\}$
correspond precisely to the states on $B$ under this isomorphism. In the special case $B=M_{n}(\mathbb{C})$ and $\tau=\operatorname{Tr}_{n}$, the canonical trace on $M_{n}(\mathbb{C})$, elements of $D_{\tau}(B)$ are called density matrices.
Example 4.18.
Let $n,m\in\mathbb{N}$, $A=M_{n}(\mathbb{C})$, $B=M_{m}(\mathbb{C})$, and $\tau$ be the canonical trace on $B$. Let $F\colon A\to B$ be a linear map and denote the standard matrix units in $M_{n}(\mathbb{C})$ by $e_{i,j}$ for $i,j=1,\ldots,n$. We recall, see for instance [Stø13, Chapter 4], that the Choi matrix $C_{F}$ of $F$ is given by
$C_{F}:=\sum_{i,j=1}^{n}e_{i,j}\otimes F(e_{i,j})\,\in\,M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})\simeq M_{nm}(\mathbb{C}).$
The assignment $F\mapsto C_{F}$ establishes an isomorphism between $\operatorname{Hom}(A,B)$ and $A\otimes B$. This isomorphism is called the Choi-Jamiolkowski isomorphism.
Moreover, we also recall that the dual functional $\widehat{F}$ on $A\otimes B=M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})$ is the linear functional satisfying
$\widehat{F}(a\otimes b)={\rm Tr}_{m}(F(a)b^{t})$
for all $a\in M_{n}(\mathbb{C})$, $b\in M_{m}(\mathbb{C})$, and that $\widehat{F}$ is positive if (and only if) $F$ is completely positive, in which case the positive density matrix corresponding to $\widehat{F}$ is the transpose of $C_{F}$, cf. [Stø13, Lemma 4.2.3 and Theorem 4.2.7]. Now, let $t$ denote the isomorphism from $B^{\rm op}=M_{m}(\mathbb{C})^{\rm op}$ onto $B=M_{m}(\mathbb{C})$ given by $t(b^{\rm op})=b^{t}$ for all $b\in B$. Then one readily verifies that
$\omega_{\tau}(F)=\widehat{F}\circ({\rm id}_{A}\otimes t).$
Thus, if $F$ is completely positive and we identify $A\otimes B^{\rm op}$ with $A\otimes B$ via the isomorphism $\operatorname{id}_{A}\otimes t$, we get that the density matrix corresponding to the positive functional $\omega_{\tau}(F)$ is the positive matrix $(C_{F})^{t}$.
Example 4.19.
Let $B$ be a finite-dimensional $C^{*}$-algebra, and $\tau$ be a faithful trace on $B$ (the existence of such a trace is guaranteed by the structure theorem for finite-dimensional $C^{*}$-algebras). In the notation of [Was24], we get a Hilbert space $(B,\langle\cdot,\cdot\rangle_{\rm KMS})$ where
$\langle x,y\rangle_{\rm KMS}=\tau(x^{*}y).$
For any linear map $F\colon B\to B$, one assigns an element $\tau^{\rm KMS}(F)\in B\otimes B^{\rm op}$ by choosing a basis $\{b_{i}\}_{i=1}^{\dim B}$ for $B$ that is orthonormal with respect to $\langle\cdot,\cdot\rangle_{\rm KMS}$ and setting
$\tau^{\rm KMS}(F)=\sum_{i=1}^{\dim B}F(b_{i})\otimes(b_{i}^{*})^{\rm op}.$
By [Was24, Proposition 3.7], $F$ is completely positive if and only if $\tau^{\rm KMS}(F)$ is positive. The functional $\varphi_{\tau^{\rm KMS}(F)}$ evaluated on a basis element $b_{m}\otimes b_{n}^{\rm op}$ of $B\otimes B^{\rm op}$ is then given by
$\varphi_{\tau^{\rm KMS}(F)}(b_{m}\otimes b_{n}^{\rm op})$ $=(\tau\otimes\tau^{\rm op})\left(\left(\sum_{i=1}^{\dim B}F(b_{i})\otimes(b_{i}^{*})^{\rm op}\right)(b_{m}\otimes b_{n}^{\rm op})\right)$
$=\sum_{i=1}^{\dim B}\tau(F(b_{i})b_{m})\tau(b_{i}^{*}b_{n})$
$=\tau(F(b_{n})b_{m})$
$=\omega_{\tau}(F)(b_{n}\otimes b_{m}^{\rm op}).$
Hence, we get that $\omega_{\tau}(F)$ and $\tau^{\rm KMS}(F)$ are related via the following equality:
$\omega_{\tau}(F)=\varphi_{\tau^{\rm KMS}(F)}\circ\Sigma^{\rm op}.$
In [CGW25] the element $\tau^{\rm KMS}(F)$ is denoted by $\operatorname{Choi}(F)$ and called the Choi matrix of $F$.
Example 4.20.
Let $A$ be a unital $C^{*}$-algebra, $m\in\mathbb{N}$, $B=M_{m}(\mathbb{C})$, and $\tau$ be the canonical trace on $B$. For each linear map $F\colon A\to M_{m}(\mathbb{C})$, let $\widehat{F}$ be the linear functional on $A\otimes M_{m}(\mathbb{C})\simeq M_{m}(A)$ determined by
$\widehat{F}(a\otimes e_{i,j})=F(a)_{i,j}\quad\text{ for every $a\in A$ and $i,j\in\{1,\ldots,m\}$},$
where the $e_{i,j}$’s denote the standard matrix units in $M_{m}(\mathbb{C})$ and $F(a)_{i,j}$ denotes the $(i,j)$-coefficient of the matrix $F(a)$. Then, as shown in [BO08, Proposition 1.5.14], the map $F$ is completely positive if and only if $\widehat{F}$ is positive, and the map $F\mapsto\widehat{F}$ gives a bijective correspondence between $\operatorname{CP}(A,M_{m}(\mathbb{C}))$ and the cone of all positive linear functionals on $A\otimes M_{m}(\mathbb{C})$. Now, it is easy to verify that $\widehat{F}(a\otimes b)={\rm Tr}_{m}(F(a)b^{t})$ for all $a\in A$ and $b\in M_{m}(\mathbb{C})$. Thus, in analogy with Example 4.18, $\widehat{F}$ might be called the dual functional of $F$ on $A\otimes M_{m}(\mathbb{C})$. Moreover, setting $t(b^{\rm op}):=b^{t}$ for all $b\in M_{m}(\mathbb{C})$, it is clear that we have $\omega_{\tau}(F)=\widehat{F}\circ({\rm id}_{A}\otimes t)$.
Remark 4.21.
In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiolkowski isomorphism in the literature. See for example [Haa21] and [HKS24] and references therein.
5 Induced metrics on trace channels
As discussed in Section 3, we will in this section determine some mild criteria, under which we can guarantee that stability and chaining hold. All results in this section are stated and proved using seminorms (rather than Lipschitz seminorms). However, the examples we have in mind arise from spectral triples through constructions explained in Section 2.2.
5.1 Stability
Since we are inducing metrics on the set of trace channels, we need to know that this set is stable under tensor products. This next result establishes this fact.
Lemma 5.1.
Let $A$ and $C$ be unital $C^{*}$-algebras, and $B$ and $D$ be $C^{*}$-algebras with faithful traces $\tau_{B}$ and $\tau_{D}$, respectively. Then for any $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ and $G\in\operatorname{TC}_{\tau_{D}}(C,D)$ we have that
$F\otimes G$ $\in\operatorname{TC}_{\tau_{B}\otimes\tau_{D}}(A\otimes_{\rm max}C,B\otimes_{\rm max}D),\text{ and }$
$F\otimes G$ $\in\operatorname{TC}_{\tau_{B}\otimes\tau_{D}}(A\otimes_{\rm min}C,B\otimes_{\rm min}D).$
Proof.
By Lemma 2.7 we have that $F\odot G$ extend to a completely positive map between both $A\otimes_{\rm max}C$ and $B\otimes_{\rm max}D$, and $A\otimes_{\rm min}C$ and $B\otimes_{\rm min}D$. Hence it remains to check that $(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=1$ in both these cases. Indeed,
$(\tau_{B}\otimes\tau_{D})((F\otimes G)(1_{A}\otimes 1_{C}))=\tau_{B}(F(1_{A}))\tau_{D}(G(1_{C}))=1.$
This completes the proof. ∎
The main technical result we need in this section is the following.
Lemma 5.2.
Let $A$ be a $C^{*}$-algebra, $B$ be a unital $C^{*}$-algebra, $\delta$ a $C^{*}$-norm on $A\odot B$, $L$ be a seminorm on $A$, and $K$ be a seminorm on $B\otimes_{\delta}A$. If the following conditions are satisfied:
  1. 1.
    $1_{B}\otimes L\leq K$, and
  2. 2.
    $K(1_{B}\otimes x)\leq 1$ for every $x\in A$ satisfying $L(x)\leq 1$.
Then
$\operatorname{mk}_{L}(\varphi_{1},\varphi_{2})=\operatorname{mk}_{K}(\psi\otimes\varphi_{1},\psi\otimes\varphi_{2}),$ (5.1)
for all $\varphi_{1},\varphi_{2}\in S(A)$ and $\psi\in S(B)$.
Proof.
We compute that for any $\varphi_{1},\varphi_{2}\in S(A)$ and $\psi\in S(B)$,
$\operatorname{mk}_{L}(\varphi_{1},\varphi_{2})$ $=\sup\{|\varphi_{1}(x)-\varphi_{2}(x)|\colon x\in A,\ L(x)\leq 1\}$
$=\sup\{|(\psi\otimes(\varphi_{1}-\varphi_{2}))(1_{n}\otimes x)|\colon x\in A,\ L(x)\leq 1\}$
$\leq\sup\{|(\psi\otimes(\varphi_{1}-\varphi_{2}))(X)|\colon X\in B\otimes_{\delta}A,\ K(X)\leq 1\}$
$=\operatorname{mk}_{K}(\psi\otimes\varphi_{1},\psi\otimes\varphi_{2}),$
where the inequality follows because $K(1_{B}\otimes x)\leq 1$ if $L(x)\leq 1$. Next we compute that,
$\operatorname{mk}_{K}(\psi\otimes\varphi_{1},\psi\otimes\varphi_{2})$ $=\sup\{|(\psi\otimes(\varphi_{1}-\varphi_{2}))(X)|\colon X\in B\otimes_{\delta}A,\ K(X)\leq 1\}$
$=\sup\{|(\varphi_{1}-\varphi_{2})((\psi\otimes\operatorname{id}_{A})(X))|\colon X\in B\otimes_{\delta}A,\ K(X)\leq 1\}$
$\leq\sup\{|\varphi_{1}(x)-\varphi_{2}(x)|\colon x\in A,\ L(x)\leq 1\}$
$=\operatorname{mk}_{L}(\varphi_{1},\varphi_{2}),$
where the inequality follows because for any $X\in B\otimes_{\delta}A$ we have that
$L((\psi\otimes\operatorname{id}_{A})(X))\leq 1_{B}\otimes L(X)\leq K(X).$
This completes the proof. ∎
We are now able to state the main theoretical result concerning stability. Note that $1_{n}^{\rm op}$ denotes the identity of $M_{n}(\mathbb{C})^{\rm op}$.
Theorem 5.3.
Let $A$ be a unital $C^{*}$-algebra, and $B$ be a $C^{*}$-algebra with a faithful trace $\tau$. For each $n\in\mathbb{N}$ suppose that we have a seminorm $L_{n}$ on $(M_{n}(\mathbb{C})\otimes A)\otimes_{\rm max}(M_{n}(\mathbb{C})\otimes B)^{\rm op}$ such that
  1. 1.
    $(1_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes L_{1})\circ\Sigma_{[23]}\leq L_{n}$, and
  2. 2.
    $L_{n}(\Sigma_{[23]}(1_{n}\otimes 1_{n}^{\rm op}\otimes x))\leq 1$ for every $x\in A\otimes_{\rm max}B^{\rm op}$ satisfying $L_{1}(x)\leq 1$.
Then the sequence $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}\}_{n\in\mathbb{N}}$, where $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$, is stable.
Proof.
To set ourselves in a situation where we can use Lemma 5.2 we define for each $n\geq 2$ a seminorm $K_{n}\colon M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}\otimes A\otimes_{\rm max}B^{\rm op}\to[0,\infty]$ by setting $K_{n}=L_{n}\circ\Sigma_{[23]}$. By conditions 1. and 2. in the hypothesis we get that
  1. 1.
    $1_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes L_{1}\leq K_{n}$, and
  2. 2.
    $K_{n}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)\leq 1$ for every $x\in A\otimes_{\rm max}B^{\rm op}$ satisfying $L_{1}(x)\leq 1$.
Furthermore, we have that for any $\varphi,\psi\in S(M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}\otimes A\otimes_{\rm max}B^{\rm op})$, the following equality holds:
$\operatorname{mk}_{K_{n}}(\varphi,\psi)=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}\varphi,\Sigma_{[23]}^{*}\psi).$ (5.2)
Now, let $F,G\in\operatorname{TC}_{\tau}(A,B)$, and $n\in\mathbb{N}$. Then
$\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm max}(\operatorname{id}_{n}\otimes F,\operatorname{id}_{n}\otimes G)$
$=\operatorname{mk}_{L_{n}}(\omega_{\operatorname{Tr}_{n}\otimes\tau}^{\rm max}(\operatorname{id}_{n}\otimes F),\omega_{\operatorname{Tr}_{n}\otimes\tau}^{\rm max}(\operatorname{id}_{n}\otimes G))$
$=\operatorname{mk}_{L_{n}}(\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F)),\Sigma_{[23]}^{*}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G)))$ by Equation 4.5,
$=\operatorname{mk}_{K_{n}}(\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(F),\omega_{\operatorname{Tr}_{n}}^{\rm max}(\operatorname{id}_{n})\otimes\omega_{\tau}^{\rm max}(G))$ by Equation 5.2,
$=\operatorname{mk}_{L_{1}}(\omega_{\tau}^{\rm max}(F),\omega_{\tau}^{\rm max}(G))$ by Lemma 5.2,
$=\Delta_{\tau,L_{1}}^{\rm max}(F,G),$
which is what we wanted to show. ∎
Remark 5.4.
If $B$ is a unital $C^{*}$-algebra with an amenable faithful trace $\tau$, then the above result and proof go through if we replace the maximal tensor product with the minimal tensor product everywhere.
5.2 Chaining
As mentioned in Section 3, we want to talk about chaining for trace channels, and it is not always the case the the composition of two trace channels is again a trace channel. We can however establish that the composition of two trace channels is again a trace channel in the following cases. If $B$ and $C$ are $C^{*}$-algebras having faithful traces $\tau_{B}$ and $\tau_{C}$, respectively, we set
$\operatorname{QC}(B,C,\tau_{B},\tau_{C}):=\{F\in\operatorname{CP}(B,C):\tau_{B}=\tau_{C}\circ F\},$
that is $\operatorname{QC}(B,C,\tau_{B},\tau_{C})$ is the set of completely positive trace-preserving maps.
Lemma 5.5.
Let $A$, $B$, and $C$ be $C^{*}$-algebras, where $A$ and $B$ are unital, and $B$ and $C$ have faithful traces $\tau_{B}$ and $\tau_{C}$, respectively. Then composition of linear maps restricts to maps
${\operatorname{UCP}(A,B)\times\operatorname{TC}_{\tau_{C}}(B,C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{TC}_{\tau_{C}}(A,C),}$${\operatorname{TC}_{\tau_{B}}(A,B)\times\operatorname{QC}(B,C,\tau_{B},\tau_{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{TC}_{\tau_{C}}(A,C).}$
Proof.
Suppose first that $F\in\operatorname{UCP}(A,B)$ and that $G\in\operatorname{TC}_{\tau_{C}}(B,C)$. It is clear that $G\circ F$ is completely positive, and we compute that
$\tau_{C}(G(F(1_{A})))=\tau_{C}(G(1_{B}))=1.$
Hence, $G\circ F\in\operatorname{TC}_{\tau_{C}}(A,C)$.
Suppose now that $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ and that $G\in\operatorname{QC}(B,C,\tau_{B},\tau_{C})$. Then we compute that
$\tau_{C}(G(F(1_{A})))=\tau_{B}(F(1_{A}))=1.$
This completes the proof. ∎
With this result in mind, we say that a pair $(F,G)$ is composable if
$(F,G)$ $\in\operatorname{UCP}(A,B)\times\operatorname{TC}_{\tau_{C}}(B,C),\text{ or }$
$(F,G)$ $\in\operatorname{TC}_{\tau_{B}}(A,B)\times\operatorname{QC}(B,C,\tau_{B},\tau_{C}).$
Remark 5.6.
Note that $\operatorname{UCP}(A,B)$ is a subset of $\operatorname{TC}_{\tau_{B}}(A,B)$ only if $\tau_{B}$ is a tracial state. This will only impact the wording of Theorem 5.10, where all traces will be states.
Before we can state the main results of this section, we are going to need some technical results. This first result says how the map $\omega_{\tau}$ interacts with compositions.
Lemma 5.7.
Let $A$, $B$, and $C$ be $C^{*}$-algebras, where $C$ has a faithful trace $\tau$. If $F\colon A\to B$ and $G\colon B\to C$ are linear, then
$\omega_{\tau}(G\circ F)=\omega_{\tau}(G)\circ(F\odot\operatorname{id}_{C^{\rm op}})$
as functionals on $A\odot C^{\rm op}$.
Proof.
For $a\in A$ and $c\in C$ we have that
$\omega_{\tau}(G\circ F)(a\otimes c^{\rm op})$ $=\tau(G(F(a))c)$
$=\omega_{\tau}(G)(F(a)\otimes c^{\rm op})$
$=(\omega_{\tau}(G)\circ(F\odot\operatorname{id}_{C^{\rm op}}))(a\otimes c^{\rm op}).$
If $L$ is a seminorm on $A\otimes_{\rm max}B^{\rm op}$, then we get a seminorm $L^{\rm op}$ on $B\otimes_{\rm max}A^{\rm op}$ by setting $L^{\rm op}=L\circ\Sigma^{\rm op}$. The following result says that this is the “correct” seminorm to consider if we want the distance between the adjoint of two channels to be the same as the distance between the original channels.
Lemma 5.8.
Let $A$ and $B$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$ and $\tau_{B}$ respectively. Let $L$ be a seminorm on $A\otimes_{\rm max}B^{\rm op}$. Suppose that $F,G\in\operatorname{TC}_{\tau_{B}}(A,B)$ are two $(\tau_{A},\tau_{B})$-adjointable maps. Then
$\Delta_{\tau_{B},L}^{\rm max}(F,G)=\Delta_{\tau_{A},L^{\rm op}}^{\rm max}(F^{\sharp},G^{\sharp}).$
If we furthermore have that $\tau_{A}$ and $\tau_{B}$ are amenable traces, and $L$ is a seminorm on $A\otimes_{\rm min}B^{\rm op}$, then
$\Delta_{\tau_{B},L}^{\rm min}(F,G)=\Delta_{\tau_{A},L^{\rm op}}^{\rm min}(F^{\sharp},G^{\sharp}).$
Proof.
Recall that $\operatorname{TC}_{\tau_{B}}(A,B)\subset\mathcal{L}_{\tau_{B}}^{\rm max}(A,B)$ by Proposition 4.6, and if $F\in\operatorname{TC}_{\tau_{B}}(A,B)$ is $(\tau_{A},\tau_{B})$-adjointable, then $F^{\sharp}\in\operatorname{TC}_{\tau_{A}}(B,A)$ by Proposition 4.17. Hence, both
$\Delta_{\tau_{B},L}^{\rm max}(F,G)$  $\quad\text{ and }\quad\Delta_{\tau_{A},L^{\rm op}}^{\rm max}(F^{\sharp},G^{\sharp})$
are well-defined quantities.
By Equation 4.8 and continuity of $\omega_{\tau_{B}}(F)$ and $\omega_{\tau_{A}}(F^{\sharp})$ in the respective maximal norms, we get that
$\omega_{\tau_{B}}^{\rm max}(F)=\omega_{\tau_{A}}^{\rm max}(F^{\sharp})\circ\Sigma^{\rm op}$
for all $x\in A\otimes_{\rm max}B^{\rm op}$. The same equality holds with $G$ in place of $F$ as well. We then compute that
$\Delta_{\tau_{B},L}^{\rm max}(F,G)$
$=\operatorname{mk}_{L}(\omega_{\tau_{B}}^{\rm max}(F),\omega_{\tau_{B}}^{\rm max}(G))$
$=\operatorname{mk}_{L}(\omega_{\tau_{B}}^{\rm max}(F^{\sharp})\circ\Sigma^{\rm op},\omega_{\tau_{B}}^{\rm max}(G^{\sharp})\circ\Sigma^{\rm op})$
$=\operatorname{mk}_{L^{\rm op}}(\omega_{\tau_{A}}^{\rm max}(F^{\sharp}),\omega_{\tau_{A}}^{\rm max}(G^{\sharp}))$
$=\Delta_{\tau_{A},L^{\rm op}}^{\rm max}(F^{\sharp},G^{\sharp}).$
In the case where $\tau_{A}$ and $\tau_{B}$ are amenable traces, the same argument given above works mutatis mutandis. ∎
Remark 5.9.
By Section 2.2: if $L_{A}$ and $L_{B}$ are seminorms on $A$ and $B$ respectively, then we have that
$(L_{A\otimes_{\rm max}B^{\rm op}})^{\rm op}=L_{B\otimes_{\rm max}A^{\rm op}}.$ (5.3)
The following Theorem and Corollary 5.12, are the main results of this section. Especially Corollary 5.12 is needed to establish that chaining holds for the class of unital completely positive maps between group $C^{*}$-algebras that arise from positive definite functions, cf. Theorem 6.9.
Theorem 5.10.
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras with faithful tracial states $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ respectively. Let $L$, $K$, and $M$ be seminorms on $A\otimes_{\rm max}B^{\rm op}$, $B\otimes_{\rm max}C^{\rm op}$, and $A\otimes_{\rm max}C^{\rm op}$ respectively. Suppose that
  1. 1.
    $F_{1},F_{2}\in\operatorname{TC}_{\tau_{B}}(A,B)$ are $(\tau_{A},\tau_{B})$-adjointable,
  2. 2.
    $G_{1},G_{2}\in\operatorname{TC}_{\tau_{C}}(B,C)$ are $(\tau_{B},\tau_{C})$-adjointable,
  3. 3.
    $(F_{1},G_{1})$, $(F_{2},G_{2})$, and $(F_{1},G_{2})$ are composable,
  4. 4.
    $K\circ(F_{1}\otimes\operatorname{id}_{C^{\rm op}})\leq M$, and that
  5. 5.
    $L^{\rm op}\circ(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})\leq M^{\rm op}$.
Then
$\Delta_{\tau_{C},M}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},K}^{\rm max}(G_{1},G_{2})+\Delta_{\tau_{B},L}^{\rm max}(F_{1},F_{2}).$
Proof.
Note that by Proposition 4.6 and Proposition 4.17, all of the quantities we will be considering here are well-defined. By the triangle inequality we have that
$\Delta_{\tau_{C},M}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},M}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{1})+\Delta_{\tau_{C},M}^{\rm max}(G_{2}\circ F_{1},G_{2}\circ F_{2}).$
We compute that the first of these terms is bounded as follows:
$\Delta_{\tau_{C},M}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{1})$
$=\operatorname{mk}_{M}(\omega_{\tau_{C}}^{\rm max}(G_{1}\circ F_{1}),\omega_{\tau_{C}}^{\rm max}(G_{2}\circ F_{1}))$
$=\operatorname{mk}_{M}((F_{1}\otimes\operatorname{id}_{C^{\rm op}})^{*}\omega_{\tau_{C}}^{\rm max}(G_{1}),(F_{1}\otimes\operatorname{id}_{C^{\rm op}})^{*}\omega_{\tau_{C}}^{\rm max}(G_{2}))$ by Lemma 5.7
$\leq\operatorname{mk}_{K}(\omega_{\tau_{C}}^{\rm max}(G_{1}),\omega_{\tau_{C}}^{\rm max}(G_{2}))$ $\text{since }K\circ(F_{1}\otimes\operatorname{id}_{C^{\rm op}})\leq M$
$=\Delta_{\tau_{C},K}^{\rm max}(G_{1},G_{2}).$
The second term is bounded by the following computation:
$\Delta_{\tau_{C},M}^{\rm max}(G_{2}\circ F_{1},G_{2}\circ F_{2})$
$=\operatorname{mk}_{M}(\omega_{\tau_{C}}^{\rm max}(G_{2}\circ F_{1}),\omega_{\tau_{C}}^{\rm max}(G_{2}\circ F_{2}))$
$=\operatorname{mk}_{M^{\rm op}}(\omega_{\tau_{A}}^{\rm max}(F_{1}^{\sharp}\circ G_{2}^{\sharp}),\omega_{\tau_{A}}^{\rm max}(F_{2}^{\sharp}\circ G_{2}^{\sharp}))$ by Lemma 5.8
$=\operatorname{mk}_{M^{\rm op}}((G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})^{*}\omega_{\tau_{A}}^{\rm max}(F_{1}^{\sharp}),(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})^{*}\omega_{\tau_{A}}^{\rm max}(F_{2}^{\sharp}))$ by Lemma 5.7
$\leq\operatorname{mk}_{L^{\rm op}}(\omega_{\tau_{A}}^{\rm max}(F_{1}^{\sharp}),(\omega_{\tau_{A}}^{\rm max}(F_{2}^{\sharp}))$ $\text{since }L^{\rm op}\circ(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})\leq M^{\rm op}$
$=\operatorname{mk}_{L}(\omega_{\tau_{B}}^{\rm max}(F_{1}),\omega_{\tau_{B}}^{\rm max}(F_{2}))$ by Lemma 5.8
$=\Delta_{\tau_{B},L}^{\rm max}(F_{1},F_{2}),$
Hence, we get that
$\Delta_{\tau_{C},M}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},K}^{\rm max}(G_{1},G_{2})+\Delta_{\tau_{B},L}^{\rm max}(F_{1},F_{2}).$
Remark 5.11.
If the traces $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ are amenable, then the above result has a version where the maximal tensor product is interchanged with the minimal tensor product instead. The proof of this is completely analogous to the one given above, so we omit the details here.
Corollary 5.12.
Let $A$, $B$, and $C$ be unital $C^{*}$-algebras with faithful traces $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ respectively, and with seminorms $L_{A}$, $L_{B}$, and $L_{C}$ respectively. Suppose that
  1. 1.
    $F_{1}\in\operatorname{UCP}(A,B)$, $F_{2}\in\operatorname{TC}_{\tau_{B}}(A,B)$ are $(\tau_{A},\tau_{B})$-adjointable,
  2. 2.
    $G_{1}\in\operatorname{TC}_{\tau_{C}}(B,C)$, $G_{2}\in\operatorname{QC}(B,C,\tau_{B},\tau_{C})$ are $(\tau_{B},\tau_{C})$-adjointable,
  3. 3.
    $L_{B}\circ F_{1}\leq L_{A}$, and that
  4. 4.
    $L_{B}\circ G_{2}^{\sharp}\leq L_{C}$.
Then
$\Delta_{\tau_{C},L_{A\otimes_{\rm max}C^{\rm op}}}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},L_{B\otimes_{\rm max}C^{\rm op}}}^{\rm max}(G_{1},G_{2})+\Delta_{\tau_{B},L_{A\otimes_{\rm max}B^{\rm op}}}^{\rm max}(F_{1},F_{2}).$
If we furthermore have that $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ are amenable, then
$\Delta_{\tau_{C},L_{A\otimes_{\rm min}C^{\rm op}}}^{\rm min}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},L_{B\otimes_{\rm min}C^{\rm op}}}^{\rm min}(G_{1},G_{2})+\Delta_{\tau_{B},L_{A\otimes_{\rm min}B^{\rm op}}}^{\rm min}(F_{1},F_{2}).$
Proof.
Note that all the pairs $(F_{1},G_{1})$, $(F_{1},G_{2})$, and $(F_{2},G_{2})$ are composable, hence we are in the setting of Theorem 5.10. By Remark 5.9 we simply need to show that the inequalities $L_{B}\circ F_{1}\leq L_{A}$ and $L_{B}\circ G_{2}^{\sharp}\leq L_{C}$ imply that
$L_{B\otimes_{\rm max}C^{\rm op}}\circ(F_{1}\otimes\operatorname{id}_{C^{\rm op}})$ $\leq L_{A\otimes_{\rm max}C^{\rm op}},\text{ and }$
$L_{B\otimes_{\rm max}A^{\rm op}}\circ(G_{2}^{\sharp}\otimes\operatorname{id}_{A^{\rm op}})$ $\leq L_{C\otimes_{\rm max}A^{\rm op}}.$
Let $X\in A\otimes_{\rm max}C^{\rm op}$. Then,
$L_{B\otimes_{\rm max}C^{\rm op}}((F_{1}\otimes\operatorname{id}_{C^{\rm op}})(X))$
$=\sup\limits_{\psi\in S(C)}L_{B}((F_{1}\otimes\psi)(X))+\sup\limits_{\varphi\in S(B)}L_{C}(((\varphi\circ F_{1})\otimes\operatorname{id}_{C^{\rm op}})(X))$
$=\sup\limits_{\psi\in S(C)}L_{B}(F_{1}((\operatorname{id}_{A}\otimes\psi)(X)))+\sup\limits_{\varphi\in S(B)}L_{C}(((F_{1}^{*}\varphi)\otimes\operatorname{id}_{C^{\rm op}})(X))$
$\leq\sup\limits_{\psi\in S(C)}L_{A}((\operatorname{id}_{A}\otimes\psi)(X))+\sup\limits_{\phi\in S(A)}L_{C}((\phi\otimes\operatorname{id}_{C^{\rm op}})(X))$
$=L_{A\otimes_{\rm max}C^{\rm op}}(X),$
where $F_{1}^{*}\varphi\in S(A)$ for every $\varphi\in S(B)$ since $F_{1}\in\operatorname{UCP}(A,B)$. Note that the proof of Proposition 4.17 specializes to show that $G_{2}^{\sharp}\in\operatorname{UCP}(C,B)$ since $G_{2}\in\operatorname{QC}(B,C,\tau_{B},\tau_{C})$. Hence, the second inequality follows by the same argument. Then, by Remark 5.9 and Theorem 5.10 we have that
$\Delta_{\tau_{C},L_{A\otimes_{\rm max}C^{\rm op}}}^{\rm max}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},L_{B\otimes_{\rm max}C^{\rm op}}}^{\rm max}(G_{1},G_{2})+\Delta_{\tau_{B},L_{A\otimes_{\rm max}B^{\rm op}}}^{\rm max}(F_{1},F_{2}).$
In the case where $\tau_{A}$, $\tau_{B}$, and $\tau_{C}$ are amenable, the inequality
$\Delta_{\tau_{C},L_{A\otimes_{\rm min}C^{\rm op}}}^{\rm min}(G_{1}\circ F_{1},G_{2}\circ F_{2})\leq\Delta_{\tau_{C},L_{B\otimes_{\rm min}C^{\rm op}}}^{\rm min}(G_{1},G_{2})+\Delta_{\tau_{B},L_{A\otimes_{\rm min}B^{\rm op}}}^{\rm min}(F_{1},F_{2})$
follows by Remark 5.11. ∎
6 Applications
6.1 The Kasparov exterior product of spectral triples
Recall that we defined unital spectral triples in Section 2.2. In this section we will show that using the Kasparov exterior product from $KK$-theory, we can from any spectral triple generate a sequence of seminorms on the matrix amplifications which naturally satisfy attractive stability properties as discussed in Section 3. This will be made precise below, see Theorem 6.2. We proceed to closely follow the exposition given in [Kaa24]. Let $(A,H_{A},\partial_{A})$ be a spectral triple, and suppose $\mathcal{B}\subseteq B(H_{B})$ be a unital *-subalgebra. We may then form the essentially self-adjoint operators
$\begin{split}\partial_{A}\otimes 1_{\mathcal{B}}&\colon{\rm Dom}(\partial_{A})\otimes H_{B}\to H_{A}\otimes H_{B}\\ 1_{\mathcal{B}}\otimes\partial_{A}&\colon H_{B}\otimes{\rm Dom}(\partial_{A})\to H_{B}\otimes H_{A}\end{split}$
where the domains are given by the algebraic tensor product and the codomains are the Hilbert space tensor products of the given Hilbert spaces. We take the closures of the operators and obtain closed, self-adjoint operators
$\begin{split}\partial_{A}\widehat{\otimes}1_{\mathcal{B}}&\colon{\rm Dom}(\partial_{A})\otimes H_{B}\to H_{A}\otimes H_{B}\\ 1_{\mathcal{B}}\widehat{\otimes}\partial_{A}&\colon H_{B}\otimes{\rm Dom}(\partial_{A})\to H_{B}\otimes H_{A}\end{split}$ (6.1)
While $\partial_{A}\widehat{\otimes}\operatorname{id}_{H_{B}}$ and $\operatorname{id}_{H_{B}}\widehat{\otimes}\partial_{B}$ do not generally have compact resolvents as self-adjoint unbounded operators on $H_{A}\otimes H_{B}$ and $H_{B}\otimes H_{A}$, respectively, we can still use them to induce metrics much in the same way as in Equation 2.2. More precisely, by identifying elements of the algebraic tensor products with operators on the Hilbert space tensor products in the obvious way, we could set
[…]
and extend by $\infty$ in accordance with Remark 2.10. Denote by $B$ the $C^{*}$-closure of $\mathcal{B}$ in $B(H_{B})$. Kaad then shows in [Kaa24, Proposition 4.10] that
$\begin{split}L_{\partial_{A}\widehat{\otimes}1_{\mathcal{B}}}=L_{\partial_{A}}\otimes 1_{B}\quad\text{and}\quad L_{1_{\mathcal{B}}\widehat{\otimes}\partial_{A}}=1_{B}\otimes L_{\partial_{A}}\end{split}$ (6.3)
as Lipschitz seminorms on $A\otimes_{\rm min}B$.
We proceed to discuss the exterior Kasparov product of unital spectral triples. Let therefore $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be unital spectral triples. We may form the exterior Kasparov product $(A\otimes_{\rm min}B,H,\partial_{A}\times\partial_{B})$, where the Hilbert space $H$, and the self-adjoint operator $\partial_{A}\times\partial_{B}$ depend on the parity of the two spectral triples as follows:
Even times even: Set $H=H_{A}\otimes H_{B}$ and endow $H$ with a $\mathbb{Z}/2\mathbb{Z}$ grading defined by
$\gamma=\gamma_{A}\otimes\gamma_{B}$
and $\partial_{A}\times\partial_{B}$ to be the closure of the symmetric self-adjoint operator
$\partial_{A}\otimes\operatorname{id}_{H_{B}}+\gamma_{A}\otimes\partial_{B}\colon{\rm Dom}(\partial_{A})\otimes{\rm Dom}(\partial_{B})\to H_{A}\otimes H_{B}$
Odd times odd: Set $H=(H_{A}\otimes H_{B})\oplus(H_{A}\otimes H_{B})$ and endow $H$ with a $\mathbb{Z}/2\mathbb{Z}$ grading defined by
$\gamma=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
The minimal tensor product $A\otimes_{\rm min}B$ is faithfully represented on $H$ via the diagonal representation
$x\mapsto\begin{pmatrix}x&0\\ 0&x\end{pmatrix}$.
Finally, consider the operator
$\begin{split}\begin{pmatrix}0&\partial_{A}\otimes\operatorname{id}_{H_{B}}+i\operatorname{id}_{H_{A}}\otimes\partial_{B}\\ \partial_{A}\otimes\operatorname{id}_{H_{B}}-i\operatorname{id}_{H_{A}}\otimes\partial_{B}&0\end{pmatrix}\end{split}$
as an operator $({\rm Dom}(\partial_{A})\otimes{\rm Dom}(\partial_{B}))^{\oplus 2}\to(H_{A}\otimes H_{B})^{\oplus 2}$. We let $\partial_{A}\times\partial_{B}$ be the closure of this operator.
Odd times even and even times odd: In both these cases $H=H_{A}\otimes H_{B}$ without any grading. If $(B,H_{B},\partial_{B},\gamma_{B})$ is even, we set $\partial_{A}\times\partial_{B}$ to be the closure of
$\partial_{A}\otimes\gamma_{B}+\operatorname{id}_{H_{A}}\otimes\partial_{B}\colon{\rm Dom}(\partial_{A})\otimes{\rm Dom}(\partial_{B})\to H_{A}\otimes H_{B}.$
If $(A,H_{A},\partial_{A},\gamma_{A})$ is even, we set $\partial_{A}\times\partial_{B}$ to be the closure of
$\partial_{A}\otimes\operatorname{id}_{H_{B}}+\gamma_{A}\otimes\partial_{B}\colon{\rm Dom}(\partial_{A})\otimes{\rm Dom}(\partial_{B})\to H_{A}\otimes H_{B}.$
We have that [Kaa24, Theorem 7.1] and its proof specialize to the following result, which we state without proof.
Theorem 6.1.
Let $(A,H_{A},\partial_{A})$ and $(B,H_{B},\partial_{B})$ be spectral triples. Then for any $x\in A\otimes_{\rm min}B$ we have that
$(1_{A}\otimes L_{\partial_{B}})(x)$ $\leq L_{\partial_{A}\times\partial_{B}}(x),\text{ and}$
$(L_{\partial_{A}}\otimes 1_{B})(x)$ $\leq L_{\partial_{A}\times\partial_{B}}(x).$
In the following, we will for ease of notation write $1$ instead of $\operatorname{id}_{H}$ for the respective identity maps appearing in the formulas for $\partial_{A}\times\partial_{B}$.
Theorem 6.2.
Let $(A,H_{A},\partial_{A})$, and $(B^{\rm op},H_{B},\partial_{B})$ be unital spectral triples, and assume that $B$ admits an amenable faithful trace $\tau$. For each $n\in\mathbb{N}$, let $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ be a spectral triple. By identifying $M_{n}(\mathbb{C})^{\rm op}$ with $M_{n}(\mathbb{C})$ we may view $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ as a spectral triple for $M_{n}(\mathbb{C})^{\rm op}$ as well. Let
$L_{1}$ $=L_{\partial_{A}\times\partial_{B}},\text{ and }$
$L_{n}$ $=L_{(\partial_{n}\times\partial_{n})\times(\partial_{A}\times\partial_{B})}\circ\Sigma_{[23]},\text{ for }n\geq 2.$
Then the sequence $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$, where $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$, is stable.
Proof.
For ease of notation, let $\partial=(\partial_{n}\times\partial_{n})\times(\partial_{A}\times\partial_{B})$. To apply Theorem 5.3 and Remark 5.4 we need to check that the following two conditions are satisfied:
  1. 1.
    $(1_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes L_{1})\circ\Sigma_{[23]}\leq L_{n}$, and
  2. 2.
    $L_{n}(\Sigma_{[23]}(1_{n}\otimes 1_{n}^{\rm op}\otimes x))\leq 1$ for every $x\in A\otimes_{\rm min}B^{\rm op}$ satisfying $L_{1}(x)\leq 1$.
By Theorem 6.1 we have that
$1_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes L_{\partial_{A}\times\partial_{B}}\leq L_{\partial}.$
Since pre-composing with $\Sigma_{[23]}$ does not change this inequality we get that
$(1_{M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op}}\otimes L_{1})\circ\Sigma_{[23]}\leq L_{n}.$
It remains to show that $L_{n}(\Sigma_{[23]}(1_{n}\otimes 1_{n}^{\rm op}\otimes x))\leq 1$ for every $x\in A\otimes_{\rm min}B^{\rm op}$ satisfying $L_{1}(x)\leq 1$. By the definition of $L_{n}$, this means that we want to show that
$L_{\partial}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)\leq 1$
for every $x\in A\otimes_{\rm min}B^{\rm op}$ satisfying $L_{\partial_{A}\times\partial_{B}}(x)\leq 1$. To prove this we need to consider the parity of the three spectral triples $(A,H_{A},\partial_{A})$, $(B^{\rm op},H_{B},\partial_{B})$, and $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$. Denote the resulting external product by
$((M_{n}(\mathbb{C})\otimes M_{n}(\mathbb{C})^{\rm op})\otimes(A\otimes_{\rm min}B^{\rm op}),H,\partial).$
The proof is a straightforward but tedious exercise in writing out Kasparov products. There are eight cases to consider, and in every case one verifies that $L_{\partial}(1_{n}\otimes 1_{n}\otimes x)=L_{\partial_{A}\times\partial_{B}}(x)$ for all $x\in A\otimes_{\rm min}B^{\rm op}$. We write out what $H$ and $\partial$ is in the following exemplary case.
Even, even, and even: In this case $H=(H_{n}\otimes H_{n})\otimes(H_{A}\otimes H_{B})$, and
$\partial=(\partial_{n}\times\partial_{n})\otimes 1+(\gamma_{n}\otimes\gamma_{n})\otimes(\partial_{A}\times\partial_{B}).$
Hence, for any $x\in A\otimes_{\rm min}B^{\rm op}$ we have that
$L_{\partial}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)$ $=\|[\partial,1_{n}\otimes 1_{n}^{\rm op}\otimes x]\|$
$=\|(\gamma_{n}\otimes\gamma_{n})\otimes[\partial_{A}\times\partial_{B},x]\|$
$=\|(\gamma_{n}\otimes\gamma_{n})\|\|[\partial_{A}\times\partial_{B},x]\|$
$=\|[\partial_{A}\times\partial_{B},x]\|$
$=L_{\partial_{A}\times\partial_{B}}(x).$
Even though $\partial$ will look more complicated in some of the other cases, it is still relatively straightforward to conclude that $L_{\partial}(1_{n}\otimes 1_{n}^{\rm op}\otimes x)=L_{\partial_{A}\times\partial_{B}}(x)$ for all $x\in A\otimes_{\rm min}B^{\rm op}$. This completes the proof. ∎
Remark 6.3.
Note that one could use two different spectral triples for $M_{n}(\mathbb{C})$ and for $M_{n}(\mathbb{C})^{\rm op}$ in Theorem 6.2. We have chosen to not do this for the ease of notation.
6.2 Group $C^{*}$-algebras and length functions
By specializing to countable discrete groups, we will in this section construct seminorms on twisted group $C^{*}$-algebras from which we can construct stable metrics on the set of trace channels, see Example 6.10. Moreover, there is a distinguished class of unital completely positive maps on twisted group $C^{*}$-algebras coming from the positive definite functions on the group, and we show that these maps behave well with respect to the chaining property for the same metrics, see Theorem 6.9.
Suppose now $G$ is a discrete group and that $\sigma\in Z^{2}(G,\mathbb{T})$ is a $2$-cocycle. The *-algebra $C_{c}(G,\sigma)$ consists of the complex functions on $G$ with finite support, equipped with convolution and involution defined by
$(f_{1}*_{\sigma}f_{2})(x)=\sum_{y\in G}f_{1}(y)f_{2}(y^{-1}x)\sigma(y,y^{-1}x)\quad\text{and}\quad f^{*}(x)=\overline{\sigma(x,x^{-1})}\overline{f(x^{-1})}$
for $f,f_{1},f_{2}\in C_{c}(G,\sigma)$ and $x\in G$. Associated with the pair $(G,\sigma)$ is also the $\sigma$-twisted left regular representation $\lambda^{\sigma}$. This is the *-representation $\lambda^{\sigma}\colon C_{c}(G,\sigma)\to B(\ell^{2}(G))$, where
$\lambda^{\sigma}(f)\xi(x)=\sum_{y\in G}f(y)\xi(y^{-1}x)\sigma(y,y^{-1}x)$
for $f\in C_{c}(G,\sigma)$, $\xi\in\ell^{2}(G)$ and $x\in G$. Thus $\lambda^{\sigma}$ is the extension of the convolution on $C_{c}(G,\sigma)$ to a bounded operator on $\ell^{2}(G)$. Taking the operator norm closure of $C_{c}(G,\sigma)$ under the (faithful) *-representation $\lambda^{\sigma}$ we obtain the $\sigma$-twisted reduced group $C^{*}$-algebra $C^{*}_{r}(G,\sigma)$. Equivalently, $C^{*}_{r}(G,\sigma)$ is the $C^{*}$-subalgebra of $B(\ell^{2}(G))$ generated $\{\lambda^{\sigma}_{g}:g\in G\}$, where $\lambda^{\sigma}_{g}:=\lambda^{\sigma}(\delta_{g})$ for each $g\in G$. Here $\delta_{g}\in\ell^{2}(G)$ is the function defined by
$\delta_{g}(x)=\begin{cases}1&x=g,\\ 0&x\neq g.\end{cases}$
Note moreover that there is a faithful canonical tracial state $\tau_{\sigma}\in S(C^{*}_{r}(G,\sigma))$ determined by
$\tau_{\sigma}(\lambda^{\sigma}(f))=f(1_{G})\quad\text{ for all }f\in C_{c}(G,\sigma).$
We may also consider the $\sigma$-twisted right regular representation $\rho^{\sigma}$. This is the *-antihomomorphism $\rho^{\sigma}\colon C_{c}(G,\sigma)\to B(\ell^{2}(G))$, where
$\rho^{\sigma}(f)\xi(x)=\sum_{y\in G}\xi(y)f(y^{-1}x)\sigma(y,y^{-1}x)$
for $f\in C_{c}(G,\sigma)$, $\xi\in\ell^{2}(G)$ and $x\in G$. This defines a faithful *-representation of the opposite algebra $C^{*}_{r}(G,\sigma)^{\rm op}$
We will want to define spectral triples for $C^{*}_{r}(G,\sigma)$, for which we record the following definition.
Definition 6.4 ([Con89]).
If $G$ is a group, a length function $l\colon G\to[0,\infty)$ is a function satisfying the following three conditions:
  1. 1.
    $l(s)=0$ if $s=1_{G}$,
  2. 2.
    $l(s^{-1})=l(s)$ for all $s\in G$, and
  3. 3.
    $l(st)\leq l(s)+l(t)$ for all $s,t\in G$.
Suppose moreover that $G$ is countable and the length function $l$ is proper, that is, for every $r\geq 0$, the set $\ell^{-1}([0,r])$ is a finite set. Such a length function gives rise to an essentially self-adjoint unbounded operator $\partial_{l}$ on $\ell^{2}(G)$ given by the linear extension of
$\begin{split}\partial_{l}\colon C_{c}(G)&\to\ell^{2}(G)\\ \delta_{s}&\mapsto l(s)\delta_{s}\end{split}$
As $\partial_{l}$ is essentially self-adjoint, we may take its self-adjoint closure, which we will also denote by $\partial_{l}$ to ease notation. We thus obtain a spectral triple $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)$ is faithfully represented on $\ell^{2}(G)$ with the left regular representation of $C_{c}(G,\sigma)$. We furthermore obtain a spectral triple $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$, where $C^{*}_{r}(G,\sigma)^{\rm op}$ is faithfully represented on $\ell^{2}(G)$ with the right regular representation of $C_{c}(G,\sigma)$. Note that a straightforward calculation such as in [Rie02, p. 614] shows that for any $f$ in $C_{c}(G,\sigma)$, the commutator $\partial_{l},\lambda^{\sigma}(f)]$ belongs to the $\sigma$-twisted crossed product $\ell^{\infty}(G)\rtimes_{{\rm red},\sigma}G$, being here identified with the $C^{*}$-subalgebra of $B(\ell^{2}(G))$ generated by $\ell^{\infty}(G)$ (acting as multiplication operators) and $\{\lambda^{\sigma}_{g}:g\in G\}$.
Let $P(G)$ denote the positive definite functions on $G$, and set
[…]
It is well-known that if $\phi\in P(G)$ (resp. $\phi\in P_{1}(G)$), then the corresponding multiplier $M_{\phi}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$, determined by
$M_{\phi}(\lambda^{\sigma}_{g})=\phi(g)\lambda^{\sigma}_{g}\quad\text{ for all }g\in G,$
is a completely positive map such that $\|M_{\phi}\|=\phi(1_{G})$ (resp. a unital completely positive map), see e.g. [BO08, Theorem 2.5.11] (when $\sigma=1$) or [BC16, Lemma 2.6].
Lemma 6.5.
Let $G$ be a discrete group, $l\colon G\to[0,\infty)$ be a length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Suppose $\phi\in P(G)$. Then $M_{\phi}$ extends to a bounded multiplier $T^{\phi}$ of the twisted crossed product $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$, and we have
$L_{\partial_{l}}(M_{\phi}(\lambda^{\sigma}(f)))\leq\|M_{\phi}\|\,L_{\partial_{l}}(\lambda^{\sigma}(f))$
for all $f\in C_{c}(G,\sigma)$. In particular, if $\phi\in P_{1}(G)$, $M_{\phi}$ is an $L_{\partial_{l}}$-seminorm contraction.
Proof.
The fact that $M_{\phi}$ extends to a completely positive map on $\ell^{\infty}(G)\rtimes_{\mathrm{red},\sigma}G$ satisfying
$T^{\phi}(\psi\lambda^{\sigma}_{g})=\phi(g)\psi\lambda^{\sigma}_{g}\quad\text{ for all }\psi\in\ell^{\infty}(G)\text{ and }g\in G,$ (6.5)
and $\|T^{\phi}\|=\|M_{\phi}\|=\phi(1_{G})$, follows by [BC15, Corollary 4.3]. Setting $\psi_{g}(x):=l(g^{-1}x)-l(x)$ for $g,x\in G$ and noting $\psi_{g}\in\ell^{\infty}(G)$ by the triangle inequality for $l$ (see [Con89, Lemma 5]), we get that
$[\partial_{l},M_{\phi}(\lambda^{\sigma}_{g})]=\phi(g)\psi_{g}\lambda^{\sigma}_{g}=T^{\phi}(\psi_{g}\lambda^{\sigma}_{g})$
for all $g\in G$. Since $\lambda^{\sigma}(f)=\sum_{g\in G}f(g)\lambda^{\sigma}_{g}$ for $f\in C_{c}(G,\sigma)$, we deduce that for every such $f$ we have
$[\partial_{l},M_{\phi}(\lambda^{\sigma}(f))]=T^{\phi}([\partial_{l},\lambda^{\sigma}(f)]).$
Hence, for each $f\in C_{c}(G,\sigma)$ we get
$L_{\partial_{l}}(M_{\phi}(\lambda^{\sigma}(f)))$ $=\|[\partial_{l},M_{\phi}(\lambda^{\sigma}(f))]\|$
$=\|T^{\phi}([\partial_{l},\lambda^{\sigma}(f)])\|$
$\leq\|T^{\phi}\|L_{\partial_{l}}(\lambda^{\sigma}(f))$
$=\|M_{\phi}\|\,L_{\partial_{l}}(\lambda^{\sigma}(f)).$
At last, if $\phi\in P_{1}(G)$, then $\|M_{\phi}\|=1$, completing the proof. ∎
Lemma 6.6.
Let $G$ be a discrete group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $\phi\in P(G)$. Then the completely positive map $M_{\phi}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, and $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$ where $\phi^{\circ}\in P(G)$ is given by
$\phi^{\circ}(g)=\phi(g^{-1})$
for all $g\in G$.
Proof.
That $\phi^{\circ}\in P(G)$ follows by [BC16, p. 7]. Hence, $M_{\phi^{\circ}}\colon C^{*}_{r}(G,\sigma)\to C^{*}_{r}(G,\sigma)$ is completely positive. For $f_{1},f_{2}\in C^{*}_{r}(G,\sigma)$ we get that
$\tau_{\sigma}(M_{\phi}(f_{1})f_{2})=(M_{\phi}(f_{1})*_{\sigma}f_{2})(1_{G})=\sum_{g\in G}\phi(g)f_{1}(g)f_{2}(g^{-1})\sigma(g,g^{-1}).$
On the other hand,
$\tau_{\sigma}(f_{1}M_{\phi^{\circ}}(f_{2}))$ $=(f_{1}*_{\sigma}M_{\phi^{\circ}}(f_{2}))(1_{G})$
$=\sum_{g\in G}f_{1}(g)\phi^{\circ}(g^{-1})f_{2}(g^{-1})\sigma(g,g^{-1})$
$=\sum_{g\in G}\phi(g)f_{1}(g)f_{2}(g^{-1})\sigma(g,g^{-1}).$
This implies that $M_{\phi}$ is $(\tau_{\sigma},\tau_{\sigma})$-adjointable, with $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$. ∎
Remark 6.7.
It is clear that $\phi(1_{G})=\phi^{\circ}(1_{G})$. Hence, $\phi\in P_{1}(G)$ if and only if $\phi^{\circ}\in P_{1}(G)$.
Proposition 6.8.
Let $G$ be a discrete group, and $\sigma\in Z^{2}(G,\mathbb{T})$. Then $P_{1}(G)$ can be identified with a subset of $\operatorname{TC}_{\tau_{\sigma}}(C^{*}_{r}(G,\sigma),C^{*}_{r}(G,\sigma))$ via the embedding $\phi\mapsto M_{\phi}$. Moreover, we have $M_{\phi}\in\operatorname{TC}_{\tau_{\sigma}}(C^{*}_{r}(G,\sigma),C^{*}_{r}(G,\sigma))$ if and only if $M_{\phi}^{\sharp}=M_{\phi^{\circ}}\in TC_{\tau_{\sigma}}(C^{*}_{r}(G,\sigma),C^{*}_{r}(G,\sigma))$.
Proof.
Let $\phi\in P_{1}(G)$. Since $M_{\phi}\in\operatorname{UCP}(C^{*}_{r}(G,\sigma),C^{*}_{r}(G,\sigma))$, we have in particular that $M_{\phi}\in\operatorname{TC}_{\tau_{\sigma}}(C^{*}_{r}(G,\sigma),C^{*}_{r}(G,\sigma))$. The result now follows by Lemma 6.6 and Remark 6.7. ∎
Theorem 6.9.
Let $G$ be a countable discrete amenable group, $l\colon G\to[0,\infty)$ be a proper length function, and $\sigma\in Z^{2}(G,\mathbb{T})$. Then the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ induced from the exterior Kasparov product of $(C^{*}_{r}(G,\sigma),\ell^{2}(G),\partial_{l})$ with $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ restricted to $P_{1}(G)$ satisfies chaining. That is
$\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}(M_{\phi_{1}}\circ M_{\phi_{2}},M_{\phi_{3}}\circ M_{\phi_{4}})\leq\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}(M_{\phi_{1}},M_{\phi_{3}})+\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}(M_{\phi_{2}},M_{\phi_{4}})$
for all $\phi_{1},\phi_{2},\phi_{3},\phi_{4}\in P_{1}(G)$.
Proof.
By [Béd95, Corollary 3.8] we have that if $G$ is amenable then $\tau_{\sigma}$ is an amenable tracial state, hence the metric $\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ is well-defined. To apply Theorem 5.10 and Remark 5.11 we only need to verify that
$L_{\partial_{l}\times\partial_{l}}\circ(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}))$ $\leq L_{\partial_{l}\times\partial_{l}},\text{ and }$
$L_{\partial_{l}\times\partial_{l}}^{\rm op}\circ(M_{\phi}^{\sharp}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}))$ $\leq L_{\partial_{l}\times\partial_{l}}^{\rm op}$
for every $\phi\in P_{1}(G)$. As in the proof of Lemma 6.5 we compute the following for $x\in C^{*}_{r}(G,\sigma)\otimes_{\rm min}C^{*}_{r}(G,\sigma)^{\rm op}$ and $\phi\in P_{1}(G)$:
$\left[\begin{pmatrix}0&\partial_{l}\otimes 1+i\otimes\partial_{l}\\ \partial_{l}\otimes 1+i\otimes\partial_{l}&0\end{pmatrix},\begin{pmatrix}(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)&0\\ 0&(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)\end{pmatrix}\right$]
$=\begin{pmatrix}0&[(\partial_{l}\otimes 1+i\otimes\partial_{l}),(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)]\\ [(\partial_{l}\otimes 1-i\otimes\partial_{l}),(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x)]&0\end{pmatrix}$
$=\begin{pmatrix}0&(T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})[(\partial_{l}\otimes 1+i\otimes\partial_{l}),x]\\ (T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})[(\partial_{l}\otimes 1-i\otimes\partial_{l}),x]&0\end{pmatrix}$
$=\begin{pmatrix}(T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})&0\\ 0&(T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})\end{pmatrix}\begin{pmatrix}0&[(\partial_{l}\otimes 1+i\otimes\partial_{l}),x]\\ [(\partial_{l}\otimes 1-i\otimes\partial_{l}),x]&0\end{pmatrix}$
Computing the operator norm of this proves that $L_{\partial_{l}\times\partial_{l}}\circ(M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}))\leq L_{\partial_{l}\times\partial_{l}}$, since
$\left\|\begin{pmatrix}(T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})&0\\ 0&(T^{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})\end{pmatrix}\right\|=1.$
For $\phi\in P_{1}(G)$ define $M_{\phi}^{\rm op}\colon C^{*}_{r}(G,\sigma)^{\rm op}\to C^{*}_{r}(G,\sigma)^{\rm op}$ by
$M_{\phi}^{\rm op}(b^{\rm op})=M_{\phi}(b)^{\rm op}.$
Note that for any $x\in C^{*}_{r}(G,\sigma)\otimes_{\rm min}C^{*}_{r}(G,\sigma)^{\rm op}$ we then have that
$\Sigma^{\rm op}((M_{\phi}\otimes\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})(x))=\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}}\otimes M_{\phi}^{\rm op}(\Sigma^{\rm op}(x)).$
We furthermore have by Lemma 6.6 that $M_{\phi}^{\sharp}=M_{\phi^{\circ}}$, Hence it suffices to show that
$L_{\partial_{l}\times\partial_{l}}\circ(\operatorname{id}_{C^{*}_{r}(G,\sigma)^{\rm op}})\otimes M_{\phi}^{\rm op})\leq L_{\partial_{l}\times\partial_{l}}$
for all $\phi\in P_{1}(G)$. This follows by the same computation as above.
With this we may appeal to Theorem 5.10 and conclude that the metric $\Delta_{\tau_{\sigma},L_{l}}^{\rm min}$ satisfies chaining when restricted to $P_{1}(G)$, which is what we wanted to show. ∎
Combining Theorem 6.2 and Theorem 6.9 gives many examples of stable metrics that satisfy chaining. We record the following example explicitly.
Example 6.10.
Let $G$ be a countable discrete amenable group, $\sigma\in Z^{2}(G,\mathbb{T})$, and $l\colon G\to[0,\infty)$ be a proper length function. As $G$ is amenable, $C^{*}_{r}(G,\sigma)$ is a nuclear $C^{*}$-algebra. This yields spectral triples $(C^{*}_{r}(G),\ell^{2}(G),\partial_{l})$ and $(C^{*}_{r}(G,\sigma)^{\rm op},\ell^{2}(G),\partial_{l})$ as we have seen above. We will obtain a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$ by choosing a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$ by Theorem 6.2. Such spectral triples have been studied in [DM21], where they show how to recover the matricial Wasserstein-1 distance from [Che+17] from the point of view of noncommutative geometry. Namely, for each $n\in\mathbb{N}$ pick $N\in\mathbb{N}$, set $H_{n}=\mathbb{C}^{n}\otimes\mathbb{C}^{N}$ and
$\partial_{n}=\sum_{i=1}^{N}L_{i}\otimes e_{i,i}$
where $L_{i}\in M_{n}(\mathbb{C})$ is a self-adjoint matrix for each $i=1,\ldots,N$. We then have a spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$ for each $n\in\mathbb{N}$. Recall that for any $\varphi\in S(M_{n}(\mathbb{C}))$, there is $\rho\in M_{n}(\mathbb{C})$ such that $\varphi=\operatorname{Tr}_{n}(\rho\,\cdot)$. D’Andrea and Martinetti show in [DM21, Proposition 5] that the associated Monge-Kantorovič metric can be realized as
$\operatorname{mk}_{\partial_{n}}(\varphi_{1},\varphi_{2})=\inf\{\operatorname{Tr}_{n}(\sqrt{u^{*}u})\mid u_{i}\in M_{n}(\mathbb{C}),\sum_{i=1}^{N}[L_{i},u_{i}]=\rho_{1}-\rho_{2}\}$
where $\varphi_{j}=\operatorname{Tr}_{n}(\rho_{j}\cdot)$. These spectral triples are therefore related to the trace norm. For each $n$, make such a choice of spectral triple $(M_{n}(\mathbb{C}),H_{n},\partial_{n})$. We then obtain a sequence of seminorms $L_{n}=L_{(\partial_{n}\times\partial_{n})\times(\partial_{l}\times\partial_{l})}\circ\Sigma_{[23]}$ by taking the appropriate Kasparov products as in Theorem 6.2, which yield a stable sequence of metrics $\{\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}\}_{n\in\mathbb{N}}$, where each $\Delta_{\operatorname{Tr}_{n}\otimes\tau,L_{n}}^{\rm min}$ is a metric on $\operatorname{TC}_{\operatorname{Tr}_{n}\otimes\tau}(M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma),M_{n}(\mathbb{C})\otimes C^{*}_{r}(G,\sigma))$. By Theorem 6.9, the metric $\Delta_{\operatorname{Tr}_{1}\otimes\tau_{\sigma},L_{1}}^{\rm min}=\Delta_{\tau_{\sigma},L_{\partial_{l}\times\partial_{l}}}^{\rm min}$ satisfies chaining with respect to the unital completely positive maps arising as multipliers $M_{\phi}$, $\phi\in P_{1}(G)$.
Data availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Conflict of interest
The authors have no conflicts of interest to declare that are relevant to the content of this article.